### v1

# ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling

Abstract
User-item interaction histories are pivotal for sequential recommendation systems but often include noise, such as unintended clicks or actions that fail to reflect genuine user preferences. To address this issue, we propose the User-Consistent Preference-based Sequential Recommendation System (ConsRec), designed to capture stable user preferences and filter noisy items from interaction histories. Specifically, ConsRec constructs a user-interacted item graph, learns item similarities from their text representations, and then extracts the maximum connected subgraph from the user-interacted item graph for denoising items. Experimental results on the Yelp and Amazon Product datasets illustrate that ConsRec achieves a 13% improvement over baseline recommendation models, showing its effectiveness in denoising user-interacted items. Further analysis reveals that the denoised interaction histories form semantically tighter clusters of user-preferred items, leading to higher relevance scores for ground-truth targets and more accurate recommendations. All codes are available at https://github.com/NEUIR/ConsRec.
ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling
Haidong Xin1thanks:   indicates equal contribution., Qiushi Xiong [^1] footnotemark: 1, Zhenghao Liu1thanks:   indicates corresponding author., Sen Mei1, Yukun Yan2, Shi Yu2, Shuo Wang2, Yu Gu1, Ge Yu1, Chenyan Xiong3 1School of Computer Science and Engineering, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3Language Technologies Institute, Carnegie Mellon University, United States
1 Introduction
Sequential recommendation systems play a pivotal role in modern digital platforms by dynamically aligning recommendations with evolving user preferences, thereby satisfying user needs and mitigating information overload Wang et al. (2019); Fang et al. (2020). While existing approaches predominantly focus on modeling user behavior through interaction histories, a critical challenge remains unaddressed: real-world interaction sequences inherently contain noise from incidental clicks and transient interests that diverge from user stable preferences. As shown in Figure 1, a user consistently purchasing Lego products might occasionally click on unrelated items like gaming consoles due to momentary curiosity. Such noisy interactions dilute the signal of genuine preferences, leading to suboptimal recommendation results.
Figure 1: Illustration of Our ConsRec Method. ConsRec constructs a user-interacted item graph to denoise items for user preference modeling.
Current methodologies primarily follow two paradigms: item-ID based modeling Kang and McAuley (2018); Sun et al. (2019) and item-content based modeling Hou et al. (2022); Liu et al. (2023). While these approaches learn representations from interaction sequences, they indiscriminately incorporate all items during training, implicitly assuming uniform relevance across interactions. This oversight amplifies noise propagation, as models struggle to distinguish between core preferences and transient outliers Lin et al. (2023b); Zhang et al. (2024a). Recent attempts to address noise either adjust item weights or rely on implicit contrastive learning Yang et al. (2023b), but these methods lack explicit mechanisms to identify and eliminate noise, limiting their interpretability.
To model the consistent preference of users, we propose the User-Consistent Preference-based Sequential Recommendation System (ConsRec), which models user stable preferences by denoising their interacted items. Specifically, ConsRec constructs a user-interacted item graph by collecting all items clicked by a specific user and connecting item pairs with a higher semantic similarity. Then, ConsRec extracts the maximum connected subgraph from the user-interacted item graph to identify the items that represent the consistent preferences of users. These identified items of the maximum connected subgraph are from the densest semantic cluster, effectively helping to model stable preferences.
Our experimental results on Amazon Product and Yelp datasets demonstrate the effectiveness of ConsRec by achieving a 13% improvement over different baseline models. ConsRec can also broaden its effectiveness to different recommendation architectures, showing its generalization ability. Further analysis shows that the retained item sets not only exhibit higher intra-cluster semantic similarity but also better match with user preferences in predicting the next item, confirming the robustness of ConsRec in broadening its effectiveness to different recommendation scenarios. Our work bridges the gap between semantic understanding and graph-based noise filtering, offering a principled solution to enhance sequential recommendation performance.
2 Related Work
Sequential recommendation systems aim to model user behaviors by leveraging user-item interaction histories. Early studies rely on the Markov Chain assumption He and McAuley (2016) and Matrix Factorization techniques Rendle (2010) for item prediction. To better represent user preferences, recent works have primarily focused on ID-based recommendation modeling, which encodes the randomly initialized ID embeddings of user-interacted items using various neural architectures. These architectures include Recurrent Neural Networks (RNNs) Hidasi et al. (2016); Liu et al. (2016); Donkers et al. (2017); Yang et al. (2017), Convolutional Neural Networks (CNNs) Tang and Wang (2018); Yuan et al. (2019); Yan et al. (2019), Graph Neural Networks (GNNs) Chang et al. (2021); Zhang et al. (2022b); Cai et al. (2023), and self-attention mechanisms Kang and McAuley (2018); Sun et al. (2019); Du et al. (2023); Yang et al. (2023a).
Instead of representing items using IDs, existing methods focus more on item content-based recommendation systems, leveraging Pretrained Language Models (PLMs) or Large Language Models (LLMs) to enhance item representations by learning textual semantics from their side information Ding et al. (2021); Yuan et al. (2023); Liu et al. (2023); Chen et al. (2024). These approaches Liu et al. (2023); Chen et al. (2024) encode the textual representations of items and further train language models using user-item interaction signals. TASTE Liu et al. (2023) verbalizes both items and user interaction histories, utilizing the T5 model Raffel et al. (2020) to align user and item representations. Leveraging the emergent abilities of LLMs Huang and Chang (2023); Lin et al. (2023a), HLLM Chen et al. (2024) introduces a hierarchical architecture for sequential recommendation comprising the item LLM and the user LLM.
User interaction sequences often contain noise, such as accidental clicks on irrelevant items Lin et al. (2023b). Existing methods focus on self-supervised denoising, which identifies noise by learning the dependencies among items from user-item interactions Lin et al. (2023b); Zhang et al. (2024a). One research direction adjusts the weight distribution of items within sequences to mitigate the influence of noise, rather than directly excluding noisy items Zhou et al. (2022); Chen et al. (2022); Yang et al. (2023b). The other direction Zhang et al. (2023, 2022a); Sun et al. (2021); Qin et al. (2021) focuses on directly identifying and removing noisy items within interaction sequences. These methods predominantly focus on item ID-based denoising approaches, which require sufficient user-item interaction signals to learn effective user and item representations Lam et al. (2008); Schein et al. (2002); Pan et al. (2019); Zhang et al. (2021). However, they overlook the significant role of text-based semantic information in identifying consistent preferences within the user-interacted item sequence Liu et al. (2023); Yuan et al. (2023).
Figure 2: The Model Architecture of ConsRec.
3 Methodology
This section introduces the User-Consistent Preference-based Sequential Recommendation System (ConsRec). We begin by presenting the preliminaries of the sequential recommendation model (Sec. 3.1). As illustrated in Figure 2, ConsRec employs a filtering mechanism to denoise user-interacted items by identifying user-consistent preferences using item similarities (Sec. 3.2) and then learning tailored item representations for similarity estimation using instruction tuning (Sec. 3.3).
3.1 Preliminary of Sequential Recommendation
The sequential recommendation task aims to predict the item $v_{t}$ that satisfies the preference of the user $\mathcal{U}$ at the $t$-th time step according to the user-item interaction history $\mathcal{H}$:
$\mathcal{H}\rightsquigarrow M_{\text{Rec}}\rightsquigarrow v_{t},$ (1)
where the user-item interaction history is represented as =$\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$. The recommendation model $M_{\text{Rec}}$ estimates the relevance between the user-item interaction history $\mathcal{H}$ and an item $v$ by calculating the ranking probability $P(v|\mathcal{H})$:
$P(v|\mathcal{H})=\text{Softmax}_{v}(\vec{\mathcal{H}}\cdot\vec{v}),$ (2)
where $\cdot$ denotes the dot product operation. Then we follow Liu et al. (2023) to get the representations $\vec{v}$ and $\vec{\mathcal{H}}$ for the candidate item $v$ and the user-item interaction history $\mathcal{H}$.
Item Representation. To fully use the semantic information of items, we verbalize the item $v$ using its identifier and $k$ associated attributes <Attr> through the following template:
$X(v)=\text{id:}v(\text{id})\dots\text{<Attr>}_{k}:v(\text{<Attr>}_{k}),$ (3)
[…]
$\vec{v}=\text{T5}(X(v)).$ (4)
User Representation. To represent the user preference, we verbalize the user-item interaction history $\mathcal{H}$ as a concatenation of the text representations (Eq. 3) of the items in $\mathcal{H}$:
$X(\mathcal{H})=X(v_{t-1});\dots;X(v_{1}),$ (5)
where ; denotes the concatenation operation. Then we encode $\mathcal{H}$ into embeddings $\vec{\mathcal{H}}$ using the T5 model, which shares the same parameters with the encoder model in Eq. 4:
$\vec{\mathcal{H}}=\text{T5}(X(\mathcal{H})).$ (6)
3.2 Denoising User-Interacted Items Using User-Consistent Preference
In user-item interactions, users may occasionally display sudden behavioral changes or lose interest in previously preferred items. Such fluctuations introduce noise into the interaction history, complicating accurate user preference modeling.
To tackle this issue, we incorporate a denoising module ($M_{\text{Filter}}$) to identify and retain items that reflect consistent user preferences. Then we train the sequential recommendation module ($M_{\text{Rec}}$) by constructing datasets using the denoised user-item interaction history $\mathcal{H}_{\text{Filter}}$ and the $t$-th step item $v_{t}$. After training, the recommendation module is used to predict the item $v_{t}$ based on $\mathcal{H}_{\text{Filter}}$:
$\mathcal{H}\rightsquigarrow M_{\text{Filter}}\xrightarrow{\mathcal{H}_{\text{Filter}}}M_{\text{Rec}}\rightsquigarrow v_{t},$ (7)
where $\mathcal{H}_{\text{Filter}}\subseteq\mathcal{H}$ represents the subset of items retained to capture consistent preferences of the user. To implement the noise filtering model $M_{\text{Filter}}$, we construct a user-interacted item graph by using the item similarity and then extract the maximum connected component from the graph to obtain the set of denoise items $\mathcal{H}_{\text{Filter}}$.
User-Interacted Item Graph. We first introduce the user-interacted item graph, which is used to distinguish some outlier items that are different from the user’s general preference.
$G_{\mathcal{U}}=(\mathcal{V},\mathcal{E}),$ (8)
where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ represents all items that have been interacted by user $\mathcal{U}$. The $\mathcal{E}$ is an edge matrix:
=$\mathcal{E}=\{e_{ij},1\leq i\leq n,1\leq j\leq n\},$ (9)
where the element $e_{ij}$ indicates whether the $i$-th node $v_{i}$ and the $j$-th node $v_{j}$ are connected with an edge. The edge is established according to the similarity score $S_{ij}$:
$e_{ij}=\begin{cases}0&\text{if }S_{ij}<\tau,\\ 1&\text{if }S_{ij}\geq\tau,\end{cases}$ (10)
where $\tau$ is the threshold of the similarity. $e_{ij}=0$ indicates that the $i$-th node $v_{i}$ and the $j$-th node $v_{j}$ are not connected, while $e_{ij}=1$ indicates that the nodes are connected. To estimate the similarity between $v_{i}$ and $v_{j}$, we encode all items in the graph node set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ into embeddings $\{\vec{v}_{1},\dots,\vec{v}_{n}\}$. Subsequently, we calculate the similarity score $S_{ij}$ using cosine similarity:
$S_{ij}=\cos(\vec{v}_{i},\vec{v}_{j}).$ (11)
Then we will introduce how to get the item embedding for similarity calculation in Sec. 3.3.
Denoising User-Interacted Items Using the Maximum Connected Component. To emphasize stronger connections among items, we aim to identify the connected components of the user-interacted item graph $G_{\mathcal{U}}$. Specifically, we perform a Breadth-First Search (BFS) to find a set of disjoint connected components:
$\mathcal{C}=\{C_{1},\dots,C_{l}\},$ (12)
where $l$ is the number of connected components, and each $C_{i}$ is a subgraph of $G_{\mathcal{U}}$. For each connected component $C_{i}$, we compute its size, defined as the number of nodes $|\mathcal{V}(C_{i})|$. The maximum connected subgraph $C_{\text{max}}$ is then selected as:
$C_{\text{max}}=\arg\max_{C_{i}}|\mathcal{V}(C_{i})|.$ (13)
These user-interacted items that belong to the node set $\mathcal{V}(C_{\text{max}})$ of the maximum connected subgraph $C_{\text{max}}$ are retained to form the denoised interaction sequence $\mathcal{H}_{\text{Filter}}$:
$\mathcal{H}_{\text{Filter}}=\{v_{j}\mid v_{j}\in\mathcal{H}\land v_{j}\in\mathcal{V}(C_{\text{max}})\}.$ (14)
This retained interaction sequence $\mathcal{H}_{\text{Filter}}$ is then passed to the recommendation module $M_{\text{Rec}}$ to predict the item at the $t$-th step. More detailed theoretical explanations are presented in Appendix A.7.
3.3 Universal Item Representation Learning for User-Interacted Graph Construction
To obtain item representations that are effective for constructing edges in the graph $G_{u}$, we adopt instruction tuning Chung et al. (2024); Wang et al. (2023) to train language models such as T5 Raffel et al. (2020) based on their textual inputs. This approach differs fundamentally from existing denoising-based methods Yuan et al. (2021a); Zhang et al. (2022a); Yang et al. (2023b); Zhang et al. (2024a), as it leverages supervised signals from different downstream recommendation tasks.
To learn text-based item representations for various recommendation tasks $\mathcal{T}$, we represent each item $v$ only through its textual semantic information. Unlike prior works relying on ID-based representations (Eq. 3), we verbalize the item only using its descriptive attributes and exclude the item ID, enhancing the generalization ability of the item representation model:
$\tilde{X}(v)^{\prime}=\text{<Attr>}_{1}:v(\text{<Attr>}_{1})\dots\text{<Attr>}_{k}:v(\text{<Attr>}_{k}).$ (15)
To fully use the user-item interaction =${\mathcal{H}}=\{v_{1},v_{2},\dots,v_{t-1}\}$, we follow previous work in recommendation model pretraining Sun et al. (2019) and use the Next Item Prediction (NIP) task to train item representations (Eq. 2). Specifically, we verbalize the user-item interaction history $\mathcal{H}$ as a concatenation of verbalized items =$\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$:
$\tilde{X}^{\prime}(\mathcal{H})=\tilde{X}(v_{t-1})^{\prime};\dots;\tilde{X}(v_{1})^{\prime},$ (16)
Then we use the instruction to illustrate the Next Item Prediction task: “This is $\mathcal{T}$ dataset. Here is the visit history list of the user: $\tilde{X}(\mathcal{H})^{\prime}$, recommend next item”, aiming to predict the $t$-th step item $v_{t}$.
4 Experimental Methodology
This section describes the datasets, evaluation metrics, baseline models, and implementation details used in our experiments.
Datasets. We utilize four sequential recommendation datasets–Beauty, Sports, Toys McAuley et al. (2015), and Yelp–to train and evaluate both the sequential recommendation module and the item filtering module. More details are shown in Appendix A.2.
Evaluation Metrics. We adopt Recall@10/20 and NDCG@10/20 to evaluate the recommendation performance, which is consistent with prior works Xie et al. (2022); Liu et al. (2023); Chen et al. (2024). Statistical significance is determined using a permutation test with P$<0.05$.
Baselines. We compare ConsRec with several widely used methods for sequential recommendation, including vanilla sequential recommendation models and some item denoising techniques.
Vanilla Sequential Recommendation Models. We compare ConsRec with both ID-based and content-based recommendation models.
First, ConsRec is evaluated against seven ID-based recommendation models. GRU4Rec Hidasi et al. (2016) employs Recurrent Neural Networks (RNNs) to model user-item interaction sequences, while SASRec Kang and McAuley (2018) and Bert4Rec Sun et al. (2019) leverage self-attention to capture user preferences from sequential interactions. Additionally, we examine models that incorporate item-side information. S3Rec Zhou et al. (2020) enhances self-attention modules through four self-supervised pretraining strategies to better model relationships among attributes, items, and users. ICAI-SR Yuan et al. (2021b) represents item-attribute relationships using a heterogeneous graph to improve relevance modeling, whereas NOVA Liu et al. (2021a) integrates item attributes as side information via attention modules. DIF-SR Xie et al. (2022) introduces a non-invasive attention mechanism to fuse item attribute information into user behavior modeling effectively.
Furthermore, ConsRec is compared with four content-based recommendation models. CoWPiRec Yang et al. (2023a) optimizes Pretrained Language Models (PLMs) to enhance item representation by aligning it with collaborative filtering representations derived from a word graph. TASTE Liu et al. (2023) leverages T5 Raffel et al. (2020) to directly encode the text representations of both users and items for matching items with user intentions. DWSRec Zhang et al. (2024b) leverages both fully whitened and relaxed whitened item representations as dual views for effective recommendations. Lastly, HLLM Chen et al. (2024) employs Large Language Models (LLMs) to generate item representations, subsequently utilizing a hierarchical architecture to predict the next item.
Item Denoising for Sequential Recommendation. We also evaluate ConsRec against three baselines that either explicitly remove noisy items or implicitly address item noise through denoising modeling. All these models are implemented based on the item-ID base recommendation model, such as Bert4Rec Sun et al. (2019). For methods that explicitly remove noisy items, DSAN Yuan et al. (2021a) learns a target item embedding and applies an adaptively sparse transformation function to filter out noisy interactions. Similarly, HSD Zhang et al. (2022a) identifies and excludes noisy interactions by detecting inconsistencies with user interests, leveraging the sequential order of recorded interactions. In contrast, DCRec Yang et al. (2023b) employs a debiased contrastive learning framework to implicitly mitigate the impact of noise.
Dataset Metrics Item ID based Recommendation Models Text based Recommendation Models
GRU4Rec Bert4Rec SASRec S3Rec NOVA ICAI-SR DIF-SR CoWPiRec DWSRec HLLM TASTE ConsRec
Beauty R@10 0.0530 0.0529 0.0828 0.0868 0.0887 0.0879 0.0908 0.0837 0.0989 0.0887 0.1030 $\text{}^{\dagger\ddagger}$
R@20 0.0839 0.0815 0.1197 0.1236 0.1237 0.1231 0.1284 0.1280 0.1496 0.1238 0.1550 $\text{}^{\dagger\ddagger}$
N@10 0.0266 0.0237 0.0371 0.0439 0.0439 0.0439 0.0446 0.0405 0.0470 0.0499 0.0517 $\text{}^{\dagger\ddagger}$
N@20 0.0344 0.0309 0.0464 0.0531 0.0527 0.0528 0.0541 0.0517 0.0597 0.0588 0.0649 $\text{}^{\dagger\ddagger}$
Sports R@10 0.0312 0.0295 0.0526 0.0517 0.0534 0.0527 0.0556 0.0588 0.0629 0.0555 0.0633 $\text{}^{\dagger\ddagger}$
R@20 0.0482 0.0465 0.0773 0.0758 0.0759 0.0762 0.0800 0.0892 0.0958 0.0816 0.0964 $\text{}^{\dagger\ddagger}$
N@10 0.0157 0.0130 0.0233 0.0249 0.0250 0.0243 0.0264 0.0300 0.0313 0.0295 0.0338 $\text{}^{\dagger\ddagger}$
N@20 0.0200 0.0173 0.0295 0.0310 0.0307 0.0302 0.0325 0.0376 0.0396 0.0360 0.0421 $\text{}^{\dagger\ddagger}$
Toys R@10 0.0370 0.0533 0.0831 0.0967 0.0978 0.0972 0.1013 0.0502 0.0967 0.0928 0.1232 $\text{}^{\dagger\ddagger}$
R@20 0.0588 0.0787 0.1168 0.1349 0.1322 0.1303 0.1382 0.0739 0.1472 0.1306 0.1789 $\text{}^{\dagger\ddagger}$
N@10 0.0184 0.0234 0.0375 0.0475 0.0480 0.0478 0.0504 0.0272 0.0457 0.0520 0.0640 $\text{}^{\dagger}$
N@20 0.0239 0.0297 0.0460 0.0571 0.0567 0.0561 0.0597 0.0331 0.0585 0.0616 0.0780 $\text{}^{\dagger}$
Yelp R@10 0.0361 0.0524 0.0650 0.0589 0.0681 0.0663 0.0698 0.0657 0.0702 0.0659 0.0738 $\text{}^{\dagger\ddagger}$
R@20 0.0592 0.0756 0.0928 0.0902 0.0964 0.0940 0.1003 0.0959 0.1103 0.0955 0.1156 $\text{}^{\dagger\ddagger}$
N@10 0.0184 0.0327 0.0401 0.0338 0.0412 0.0400 0.0419 0.0416 0.0382 0.0407 0.0397 $\text{}^{\ddagger}$
N@20 0.0243 0.0385 0.0471 0.0416 0.0483 0.0470 0.0496 0.0497 0.0482 0.0504 0.0502 $\text{}^{\dagger\ddagger}$
[…]
Implementation Details. Our ConsRec is implemented using OpenMatch Liu et al. (2021b); Yu et al. (2023) and optimized using in-batch negatives. Both the item representation model in the filter module and the recommendation module are initialized from the T5-base checkpoint [^1] https://huggingface.co/google-t5/t5-base in Huggingface Transformers Wolf et al. (2019). The item representation model in the filter module is trained using the Adam optimizer with a learning rate of 1e-4 and a batch size of 32. For the recommendation module, we follow the settings in Liu et al. (2023), which is trained using the Adam optimizer with a learning rate of 1e-4, a warm-up proportion of 0.1, and a batch size of 8. We set the filtering threshold $\tau$ to 0.7, based on the hyperparameter sensitivity analysis reported in Appendix A.4. During inference, we excluded the last interaction from the user-interaction sequence when constructing the consistency graph, thereby avoiding potential data leakage.
5 Experimental Results
In this section, we begin by evaluating the recommendation performance of ConsRec and perform ablation studies to assess the effectiveness of denoising modules and various training strategies. Then, we analyze text similarities among retained items to evaluate the effectiveness of the item representation learning method. Finally, case studies are shown in Appendix A.6.
5.1 Overall Performance
In this section, we first implement ConsRec based on TASTE Liu et al. (2023) to evaluate its recommendation performance. Subsequently, we implement ConsRec with Bert4Rec Sun et al. (2019) to evaluate its denoising capability by comparing it against various baselines.
Recommendation Performance. As shown in Table 1, we present the overall performance of ConsRec and various recommendation baseline models. Overall, ConsRec demonstrates its effectiveness by significantly outperforming all baselines across four datasets, leveraging TASTE Liu et al. (2023) as its backbone architecture.
Among all baseline models, text-based recommendation models generally exhibit significantly stronger performance compared to item ID-based models. This superiority derives from the ability of PLMs and LLMs to effectively capture matching signals through text representations of users and items. Based on the TASTE backbone, ConsRec integrates a denoising module to remove noisy items, utilizing these retained user-item interactions to train the recommendation module. The evaluation results reveal a significant improvement in the recommendation performance, highlighting its effectiveness in enabling the recommendation module to better model user behavior. Among various recommendation tasks, ConsRec demonstrates much better performance on the Beauty and Sports datasets, while showing relatively weaker improvements on the Yelp dataset. This discrepancy may be attributed to the tendency of users to make more casual clicks driven by instant interest in these shopping scenarios.
Dataset Method R@10 R@20 N@10 N@20
Beauty DSAN 0.0215 0.0320 0.0103 0.0132
HSD 0.0423 0.0645 0.0218 0.0262
DCRec 0.0609 0.0869 0.0317 0.0387
ConsRec 0.0627 0.0904 0.0336 0.0406
Sports DSAN 0.0107 0.0159 0.0051 0.0065
HSD 0.0214 0.0317 0.0103 0.0133
DCRec 0.0262 0.0407 0.0139 0.0175
ConsRec 0.0275 0.0428 0.0147 0.0185
Toys DSAN 0.0172 0.0258 0.0086 0.0104
HSD 0.0344 0.0517 0.0172 0.0209
DCRec 0.0443 0.0660 0.0243 0.0296
ConsRec 0.0468 0.0691 0.0258 0.0314
Yelp DSAN 0.0153 0.0218 0.0084 0.0116
HSD 0.0326 0.0483 0.0165 0.0247
DCRec 0.0387 0.0562 0.0204 0.0262
ConsRec 0.0398 0.0576 0.0211 0.0274
Table 2: Denoising Performance of Different Methods. To ensure a fair comparison, all denoising methods are implemented using Bert4Rec Sun et al. (2019).
Denoising Effectiveness. We compare ConsRec with three state-of-the-art denoising methods–DSAN Yuan et al. (2021a), HSD Zhang et al. (2022a), and DCRec Yang et al. (2023b)–to evaluate its denoising effectiveness. The baselines are categorized into two groups–explicit denoising methods (DSAN and HSD) and implicit denoising methods (DCRec). All methods are implemented using the item ID-based recommendation model, Bert4Rec Sun et al. (2019). The evaluation results are shown in Table 2.
Among the baseline methods, DCRec achieves the best performance by modeling user behaviors based on raw user-item interactions and effectively mitigating noise through dynamically assigning weights to user-interacted items. However, this implicit denoising approach necessitates model-specific training, restricting our experiments to the Bert4Rec model used in constructing the DCRec framework. Benefiting from explicit denoising methods, ConsRec can be applied to the item ID-based recommendation model and achieves a 4.7% improvement over all denoising baseline models, highlighting its effectiveness in filtering noisy items. This consistent improvement with the item ID-based model underscores the potential of extending the filtering approach of ConsRec to different architectures of recommendation models.
Dataset Method R@20 N@20
Beauty ConsRec (Zero-Shot) 0.1631 0.0697
w/o $M_{\text{Filter}}$ 0.1550 0.0649
[…]
w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
Sports ConsRec (Zero-Shot) 0.1014 0.0420
w/o $M_{\text{Filter}}$ 0.0964 0.0421
[…]
w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
Toys ConsRec (Zero-Shot) 0.1797 0.0786
w/o $M_{\text{Filter}}$ 0.1789 0.0780
[…]
w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
Yelp ConsRec (Zero-Shot) 0.1158 0.0504
w/o $M_{\text{Filter}}$ 0.1156 0.0502
[…]
w/o $M_{\text{Rec}}$ Training 0.1165 0.0504
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
[…]
5.2 Ablation Study
This section presents an ablation study to evaluate the contribution of the denoising module ($M_{\text{Filter}}$) and the effectiveness of our proposed training strategy in ConsRec. More experimental results are shown in Appendix A.5.
Table 3 shows the performance of several ablated versions of our model across four datasets. Comparing ConsRec (w/o $M_{\text{Filter}}$) with ConsRec (Zero-Shot), we observe a consistent improvement in recommendation performance across all datasets when a zero-shot $M_{\text{Filter}}$ is applied during inference. This highlights the inherent benefit of even an untrained denoising module in refining user-interacted items for recommendation. The impact of the denoising module becomes more evident when comparing ConsRec (Zero-Shot) with ConsRec (w/o $M_{\text{Rec}}$ Training). Training the $M_{\text{Filter}}$ module using our instruction-based method leads to a significant improvement, demonstrating the effectiveness of learning tailored item representations through denoising. Specifically, ConsRec (w/o $M_{\text{Rec}}$ Training) consistently outperforms ConsRec (w/o $M_{\text{Filter}}$), indicating that a well-trained filter is crucial for recommendation systems.
Furthermore, the full ConsRec model, which trains the recommendation module $M_{\text{Rec}}$ on the retained interactions, achieves the best performance across all datasets, with an average improvement of over 12% compared to ConsRec (Zero-Shot). This substantial gain underscores the importance of aligning the recommendation module with the retained data, enabling it to better capture user preferences and behaviors from the refined interaction history. The consistent statistically significant improvements of ConsRec over both ConsRec (w/o $M_{\text{Filter}}$) and ConsRec (Zero-Shot) further validate the effectiveness of our complete model.
(a) Average Similarity Scores among $M_{\text{Filter}}$-Retained Items.
(b) Similarity Scores between $M_{\text{Filter}}$-Retained Items and the Target Items.
Figure 3: Effectiveness of ConsRec in Learning Item Representations for Denoising. Figure 3(a) and Figure 3(b) illustrate the item similarities among $M_{\text{Filter}}$-retained items and the target items.
(a) Raw User-Item Interaction Trained $M_{\text{Rec}}$.
(b) $M_{\text{Filter}}$-Retained Items Trained $M_{\text{Rec}}$.
Figure 4: Performance of Recommendation Models Using Different Item Denoising Methods.
5.3 Effectiveness of Different Item Representation Strategies in Denoising
To assess the effectiveness of the denoising module in ConsRec, we conduct a comparative analysis involving various item representation strategies. Specifically, we compare the full ConsRec model with its variant without the denoising module (ConsRec w/o Filter), as well as versions of ConsRec equipped with different item representation methods within the denoising module. We implement the denoising module $M_{\text{Filter}}$ using three approaches: an ID-based recommendation model, Bert4Rec Sun et al. (2019), and a dense retriever, BGE-Base Xiao et al. (2023), which serve as baselines for evaluating representation effectiveness.
As illustrated in Figure 3, we analyze the text similarity among the items retained by $M_{\text{Filter}}$ using the BGE-Base model to compute pairwise similarity scores. In Figure 3(a), we report the average similarity among the retained items using different denoising methods. The results show that all methods improve the internal similarity of the retained items. Notably, when using ConsRec, the retained user-interacted items achieve the highest average similarity, indicating stronger coherence among these retained items. Furthermore, Figure 3(b) presents the average similarity between the $M_{\text{Filter}}$-retained user-interacted items and the ground-truth target item. Among all methods, ConsRec yields the highest similarity, demonstrating its effectiveness in retaining items that are semantically aligned with the target item that users needed.
As illustrated in Figure 4, we evaluate the effectiveness of various denoising methods across different recommendation systems. Figures 4(a) and 4(b) report the performance of the recommendation module $M_{\text{Rec}}$ when trained on either the raw user-item interaction set $\mathcal{H}$ or the denoised subset $\mathcal{H}_{\text{Filter}}$ retained by $M_{\text{Filter}}$. In both cases, ConsRec consistently outperforms other item representation models, with particularly strong gains at larger ranking cutoffs. This demonstrates its effectiveness in helping the recommendation model prioritize more relevant items, thereby better satisfying user preferences. The advantage of ConsRec can be attributed to its PLM-based ability to comprehend item text semantics and to capture item similarity patterns from user-item interaction signals. After training the recommendation module $M_{\text{Rec}}$ using denoised user-item interactions, ConsRec achieves substantial performance improvements, further highlighting its effectiveness.
5.4 The Effectiveness of Item Denoising in Enhancing the Recommendation Module
(a) TASTE.
(b) ConsRec.
Figure 5: Embedding Visualization of Items in Beauty. We randomly select a user and then use t-SNE to visualize item embeddings using TASTE and ConsRec.
As illustrated in Figure 5, we investigate how the retained items contribute to improving the recommendation module during training. Specifically, we get the item representations retained and removed by $M_{\text{Filter}}$ using TASTE and ConsRec.
Figure 5(a) shows the embedding space learned by TASTE, which is trained without item denoising. In this setting, the model yields a homogeneous embedding space, where retained and removed items are mixed together without clear separation. In contrast, Figure 5(b) demonstrates that ConsRec is able to better distinguish the removed items by mapping them to the periphery of the embedding space, effectively treating them as outliers. This observation highlights the effectiveness of ConsRec in leveraging denoising signals as prior knowledge to guide the recommendation module. By doing so, it facilitates a more discriminative representation space, where retained and removed items are clearly separated through contrastive training.
6 Conclusion
This paper introduces the User-Consistent Preference based Sequential Recommendation System (ConsRec), a novel framework designed to model user-consistent preferences for denoising user-interacted items. By explicitly filtering out noisy items, ConsRec enhances the interpretability of denoising and demonstrates its generalization ability, extending its advantages across various recommendation models.
Limitation
While ConsRec demonstrates effective denoising capabilities for sequential recommendations, the approach has several limitations that must be acknowledged. First, the performance of ConsRec is contingent on the quality of item descriptions; incomplete or low-quality descriptions may hinder the model’s ability to effectively filter out noisy user-item interactions. Secondly, for cold-start users with limited interaction history, ConsRec needs to build new user-item graphs from scratch, potentially introducing latency in generating personalized recommendations. This limitation could be mitigated by employing incremental graph updates, which dynamically add new edges as interactions occur, thereby reducing the need for full graph recomputation.
Appendix A Appendix
A.1 License
We show the licenses of the datasets that we use. Yelp [^2] https://www.yelp.com/dataset uses Apache License 2.0, while Amazon Product shows its terms of use at website [^3] http://jmcauley.ucsd.edu/data/amazon/. All of these licenses and agreements allow their data for academic use.
A.2 Additional Experimental Details of Data Preprocessing
Dataset Data Information Split
#Users #Items #Actions Train Dev Test
Beauty 22,363 12,101 198,502 131,413 22,363 22,363
Sports 35,598 18,357 296,337 189,543 35,598 35,598
Toys 19,412 11,924 167,597 109,361 19,412 19,412
Yelp 30,499 20,068 317,182 225,685 30,499 30,499
Table 4: Statistics of the Datasets Used for Modeling the Sequential Recommendation Module ($M_{\text{Rec}}$).
Dataset Data Information Split
#Users #Items #Actions Train Dev
Beauty 22,363 12,101 198,502 225,685 19,412
Sports 35,598 18,357 296,337 225,685 19,412
Toys 19,142 11,924 167,597 225,685 19,412
Yelp 30,499 20,068 317,182 225,685 19,412
Total 107,602 62,450 979,618 902,740 77,648
Table 5: Statistics of Training Data Used in the Filter Module ($M_{\text{Filter}}$). The data is used for optimizing item representations to estimate their similarities.
In our experiments, we employ Recbole Zhao et al. (2021) for preprocessing the data for sequential recommendation modeling, maintaining the same experimental settings as prior works Zhou et al. (2020); Xie et al. (2022); Liu et al. (2023); Chen et al. (2024). Specifically, each user-item interaction sequence $\{v_{1},\dots,v_{t}\}$ is processed using the leave-one-out strategy Xie et al. (2022); Chen et al. (2018); Sun et al. (2019); Zhou et al. (2020) to construct the training, development, and testing sets. In detail, the testing set predicts the $t$-th item $v_{t}$ based on the previously interacted items $v_{1:t-1}$ of the user, while the development set predicts the $t-1$-th item $v_{t-1}$ given the interactions $v_{1:t-2}$. The training set predicts $v_{i}$ using $v_{1:i-1}$, where $11<i<t-$ 1. Items and users with fewer than five interactions are filtered out, and all user-item interactions are treated as implicit feedback Chen et al. (2018); Sun et al. (2019); Zhou et al. (2020); Xie et al. (2022); Liu et al. (2023); Chen et al. (2024). The data statistics are shown in Table 4.
Then, we construct the dataset for training the language model to represent items for estimating their similarities (Eq. 11). We start with the training and development sets used in sequential recommendation modeling and further process them for the Next Item Prediction (NIP) task. We first resample user-item interaction sequences from different datasets, namely Beauty, Sports, Toys, and Yelp. Then, we balance the number of data points to construct the training and development sets. This approach mitigates the risk of overfitting caused by data imbalance of different recommendation tasks. As shown in Table 5, the data statistics for both the training and development sets used in item representation learning are presented.
Following previous studies Xie et al. (2022); Liu et al. (2023), we conduct evaluations in the full ranking testing scenario Dallmann et al. (2021); Krichene and Rendle (2020), where items are ranked across the entire item set rather than sampling subsets. This evaluation set mitigates the inconsistencies often observed in sampled evaluations, offering a more realistic and comprehensive assessment of recommendation performance.
A.3 Additional Details of Baseline Models
We reproduce all item-ID based models and DWSRec Zhang et al. (2024b) using implementations provided by RecBole Zhao et al. (2021), maintaining consistent settings across all models for fair comparison. For CoWPiRec and TASTE, we adhere to the default experimental settings as specified in their original papers. For HLLM, due to computational constraints, we ultilize TinyLlama-1.1B Zhang et al. (2024c) as the item-LLM component, and the user-LLM is implemented as a custom Llama model Touvron et al. (2023) with two hidden layers, trained from scratch using the following settings: learning rate of 1e-4, batch size of 2, training for 5 epochs, and a negative sampling ratio of 1024 negative samples per positive sample.
(a) SASRec
(b) BERT4Rec
(c) TASTE
Figure 6: Hyperparameter Sensitivity Analysis. We conduct experiments on the threshold hyperparameter, $\tau$, with results derived from the Amazon Beauty dataset using the SASRec 6(a), BERT4Rec 6(b), and TASTE 6(c) backbone models.
A.4 Hyperparameter Sensitivity Analysis
We investigate the sensitivity of the key hyperparameter $\tau$, which determines the similarity threshold during user-item interaction graph construction, as described in Sec. 3.2. We systematically vary $\tau$ from 0.2 to 0.8 while maintaining all other experimental settings. Figure 6 illustrates the performance across four evaluation metrics (Recall@10/20 and NDCG@10/20) for three backbone models: SASRec, Bert4Rec, and TASTE.
The results reveal a consistent pattern across all models and metrics. At lower $\tau$ values (0.2-0.4), the filtering mechanism is overly permissive, allowing to much noise in the interaction data and resulting in suboptimal performance. As $\tau$ increases to moderate values (0.5-0.6), we observe substantial performance improvements as irrelevant interactions are effectively filtered out. For instance, SASRec achieves peak R@10 performance at approximately $\tau=0.6$, with similar bell-shaped trends evident across other models and metrics. When $\tau$ exceeds 0.7, the filtering becomes excessively strict, potentially discarding valuable interaction signals and causing performance deterioration.
A.5 Ablation Study on Different Filter Models
To evaluate the effectiveness of the $M_{\text{Filter}}$ component in ConsRec, we conduct ablation studies comparing our approach with two alternative filter models–a vanilla T5 $M_{\text{Filter}}$ initialized from the T5-base checkpoint without additional training, and a TASTE $M_{\text{Filter}}$ trained specifically on recommendation data.
Dataset Method R@20 N@20
Beauty ConsRec 0.1774 0.0772
w/ T5 $M_{\text{Filter}}$ 0.1761 0.0762
w/ TASTE $M_{\text{Filter}}$ 0.1767 0.0767
Sports ConsRec 0.1082 0.0496
w/ T5 $M_{\text{Filter}}$ 0.1073 0.0457
w/ TASTE $M_{\text{Filter}}$ 0.1077 0.0471
Toys ConsRec 0.1896 0.0826
w/ T5 $M_{\text{Filter}}$ 0.1879 0.0802
w/ TASTE $M_{\text{Filter}}$ 0.1883 0.0811
Yelp ConsRec 0.1173 0.0509
w/ T5 $M_{\text{Filter}}$ 0.1169 0.0507
w/ TASTE $M_{\text{Filter}}$ 0.1170 0.0508
Table 6: Ablation Study on Different Filter Models in ConsRec.
As shown in Table 6, ConsRec consistently outperforms both alternative implementations across all datasets, demonstrating the effectiveness of our proposed training strategy for the filter module. The vanilla T5 $M_{\text{Filter}}$ exhibits the weakest performance, as it lacks domain-specific knowledge for recommendation tasks. The TASTE $M_{\text{Filter}}$ shows improved performance due to its training on recommendation tasks, but its effectiveness is limited by exposure to only a single dataset during training. In contrast, our ConsRec $M_{\text{Filter}}$ benefits from data-balanced pre-training across multiple datasets, resulting in more robust and generalizable item representations. This confirms that our training approach effectively enhances the model’s ability to capture relevant item characteristics while maintaining strong generalization capabilities across different recommendation scenarios.
Case #1 in Amazon Beauty
Node_ID: $n_{1}$    Name: Bare Escentuals bareMinerals Purely Nourishing Moisturizer: Combination Skin
Description:Hydrate skin with this lightweight, remarkably effective moisturizer for improved firmness, elasticity …
[…]
Description: … An effective and groundbreaking anti-aging moisturizer contains advanced hydration complexity that deliver …
[…]
Description: … Valentina By Valentino 1.7 Oz Eau De Parfum Spray
[…]
Description: … An oil-free nightly moisturizer that provides weightless hydration, combats excess oil and refines the look of …
Similarity: $S_{1,2}=0.4441$, $S_{1,3}=0.2527$, $S_{1,4}=0.4608$, $S_{2,3}=0.2125$, $S_{2,4}=0.4477$, $S_{3,4}=0.2049$
Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
The maximum connected subgraph $C_{max}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Filter}}$ is: $\left\{1,2,4\right\}$
Case #2 in Yelp
Node_ID: $n_{1}$    Address: Las Vegas, NV    Name: Maggiano’s Little Italy
Description:Nightlife, Restaurants, Italian, Bars …
[…]
Description:Nightlife, Food, Beer, Wine & Spirits, Pizza, Burgers …
[…]
Description:Veterinarians, Pets, Pet Services
[…]
Description:Restaurants, Beer, Wine & Spirits, Food, Vegetarian, …, Seafood, Bars, Nightlife
Similarity: $S_{1,2}=0.5926$, $S_{1,3}=0.4699$, $S_{1,4}=0.5121$, $S_{2,3}=0.4865$, $S_{2,4}=0.5150$, $S_{3,4}=0.3932$
Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
The maximum connected subgraph $C_{max}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Filter}}$ is: $\left\{1,2,4\right\}$
Table 7: Case Studies. We present cases from two datasets, Amazon Beauty and Yelp. Different colors are highlighted to distinguish preference types: Blue represents majority preference, and Orange represents discrete interaction.
A.6 Case Studies
In Table 7, we present two examples from the Amazon Beauty and Yelp datasets to evaluate the effectiveness of ConsRec. For each dataset, a user is randomly selected, and we illustrate the process of denoising user-interacted items by modeling user-consistent preferences. In the case studies, we detail the procedure of identifying the maximum connected subgraph from the user-interacted item graph for item denoising.
As shown in the first case from the Amazon Beauty dataset, the user primarily interacts with face cream products focused on moisturizing, such as “Bare Escentuals bareMinerals Purely Nourishing Moisturizer” and “Ole Henriksen Truth Creme Advanced Hydration”. However, the user accidentally clicks on a product–“VALENTINO VALENTINA by Valentino for WOMEN: EAU DE PARFUM SPRAY 1.7 OZ”–due to their instant interests, which are significantly different in terms of both product types and attributes. In detail, while other items in the sequence emphasize “hydrating” and “moisturizing”, the “Parfum Spray” is unrelated to these preferences. When ConsRec constructs the user-interacted item graph, the “Parfum Spray” product fails to form edges with other items, indicating that this interaction is likely an outlier and should be excluded from user-interacted items. Consequently, the denoised item sequence $\mathcal{H}_{\text{Filter}}$ includes only products aligned with the consistent preferences of a specific user, such as moisturizing creams, enabling the recommendation module to better model user behavior from the user-item interactions.
In the second case, we present a specific user from the Yelp dataset who visits various locations. This user primarily interacts with places characterized by attributes such as “Nightlife”, “Restaurants”, and “Bars”, demonstrating a consistent pattern of preferences. However, one of the interacted locations– the “Spay & Neuter Center of Southern Nevada”–is a pet hospital, which is distinct from the user-consistent preference for nightlife and dining. Although this pet hospital is situated on the same street as other frequently visited locations, it likely represents an occasional visit rather than a reflection of the general behavior of the user. ConsRec assigns a lower similarity score between this noisy item and other user-interacted items, effectively excluding the noisy item from the sequence of user-interacted history. This enables the recommendation model to focus on more representative items, such as “Maggiano’s Little Italy” and “Weera Thai Restaurant”.
A.7 Theoretical Analysis of Consistent User Preference Modeling via ConsRec
Preliminary. Let =$\mathcal{H}=\{v_{1},\dots,v_{n}\}$ be the full interaction history of a user. The consistent user preference set $P$ can be represented:
$P\subseteq\mathcal{H},\lvert P\rvert=m,$ (17)
and the remaining items set can be denoted:
$N=\mathcal{H}\setminus P,\lvert N\rvert=n-m.$ (18)
Then, we denote the user interaction graph as an undirected graph:
$G_{\mathcal{U}}=(\mathcal{H},\mathcal{E}).$ (19)
Here $(v_{i},v_{j})\in\mathcal{E}\Longleftrightarrow\cos(\vec{v}_{i},\vec{v}_{j}\bigr{)}=S_{ij}\geq\tau$, where the fixed threshold $\tau\in(0,1)$. In the user interaction graph $G_{\mathcal{U}}$, if every pair of nodes in a set can be connected by a path, then this set of nodes is called a “connected component”. We denote the collection of all connected components as:
$\mathcal{C}=\{C_{1},C_{2},\dots,C_{\ell}\},$ (20)
where each connected component $C_{k}\subseteq\mathcal{H}$ and $C_{i}\cap C_{j}=\emptyset$ for all $i\neq j$. Then, we denote
$C_{\text{max}}=\mathrm{arg}\max_{1\leq k\leq\ell}\lvert C_{k}\rvert$ (21)
as the connected component with the maximum number of nodes. Above all, our main claim is that the maximum connected component $C_{\text{max}}$ is exactly the user preference set $P$.
Assumptions. We need the following three weakest conditions to support the subsequent derivation.
(i) Every pair of nodes belongs to $P$ is similar:
$\forall v_{i},v_{j}\in P:S_{ij}\geq\tau.$ (22)
(ii) There is no edge between any noise node in $N$ and any preferred node in $P$:
$\forall v_{i}\in P,v_{j}\in N:S_{ij}<\tau;$ (23)
(iii) Even if noise nodes have high similarity and form small clusters, their scale will not exceed that of the preference set $P$:
$\forall\{C_{1}^{N},\ldots,C_{q}^{N}\}\subseteq\mathcal{C}:\max_{1\leq r\leq q}\lvert C_{r}^{N}\rvert\mathrm{<}m.$ (24)
Lemmas. The following three lemmas are correspond to assumptions (i)-(iii), respectively.
(1) The preference set is the connected component. By assumption (i), $\forall v_{i},v_{j}\in P$, then $S_{ij}\geq\tau$, so that $(v_{i},v_{j})\in\mathcal{E}$. If any two nodes in the set $P$ are directly connected to each other, then $P$ forms a complete graph, which necessarily constitutes a connected component.
(2) Noise nodes and preference nodes are not in the same component. By assumption (ii), $\forall v_{i}\in P,v_{j}\in N$, then $S_{ij}<\tau$, which implies that there is no direct edge between $v_{i}$ and $v_{j}$. If a noise node and a preferred node are in the same connected component, there must exist a path between the two nodes. However, as long as there is one transition from a noise node to a preferred node, it will violate assumption (ii). Therefore, such a path cannot exist. Consequently, all noise nodes and the connected component belong to different components.
(3) There is an upper bound on the scale of the noise cluster. By assumption (iii), any connected component $C_{r}^{N}$ composed solely of noise nodes satisfies $\rvert C_{r}^{N}\rvert\mathrm{<}m$.
Proof of Proposition. We prove the two directions $v\in P\iff v\in C_{\max}$.
If $v\in P$, then $v\in C_{\max}$. By lemma (1) and lemma (3), any two nodes in $P$ are connected by an edge, so that $P$ itself forms a connected component, which we denote by $C_{P}$, and $\lvert C_{P}\rvert=\lvert P\rvert=m$. Then, each noise node has no edges with $P$, and there are no edges between noise nodes. So all noise nodes are either isolated or form small components with size less than $m$. Therefore, among all components, $\lvert C_{P}\rvert$ is the largest. Hence $C_{\max}=C_{P}$, and thus $P\subseteq C_{\text{max}}$.
If $v\in C_{\max}$, then $v\in P$. Suppose, for contradiction, there exists $v_{k}\in N\cap C_{\max}$. Since $C_{\max}=C_{P}$, $v_{k}$ is connected to some $v_{i}\in P$, and there must be a path passing through several edges. However, by lemma (2) and lemma (3), there is no noise node connected to any node in $P$, and there are no edges between noise nodes. So, it is impossible to construct a path from $v_{k}$ to $v_{i}$. Therefore, noise nodes cannot appear in $C_{\max}$. Hence, $C_{\max}\subseteq P$.
Thus, we can prove that $C_{\text{max}}=P$.

### v2

@@
-ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling
-Abstract
+LISRec: Modeling User Preferences with Learned Item Shortcuts for Sequential Recommendation
+Abstract.
 User-item interaction histories are pivotal for sequential recommendation systems but often include noise, such as unintended clicks or actions that fail to reflect genuine user preferences.
-To address this issue, we propose the User-Consistent Preference-based Sequential Recommendation System (ConsRec), designed to capture stable user preferences and filter noisy items from interaction histories.
-Specifically, ConsRec constructs a user-interacted item graph, learns item similarities from their text representations, and then extracts the maximum connected subgraph from the user-interacted item graph for denoising items.
-Experimental results on the Yelp and Amazon Product datasets illustrate that ConsRec achieves a 13% improvement over baseline recommendation models, showing its effectiveness in denoising user-interacted items.
-Further analysis reveals that the denoised interaction histories form semantically tighter clusters of user-preferred items, leading to higher relevance scores for ground-truth targets and more accurate recommendations.
-All codes are available at https://github.com/NEUIR/ConsRec.
-ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling
-Haidong Xin1thanks:   indicates equal contribution., Qiushi Xiong [^1] footnotemark: 1, Zhenghao Liu1thanks:   indicates corresponding author., Sen Mei1, Yukun Yan2, Shi Yu2, Shuo Wang2, Yu Gu1, Ge Yu1, Chenyan Xiong3 1School of Computer Science and Engineering, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3Language Technologies Institute, Carnegie Mellon University, United States
-1 Introduction
-Sequential recommendation systems play a pivotal role in modern digital platforms by dynamically aligning recommendations with evolving user preferences, thereby satisfying user needs and mitigating information overload Wang et al.
-(2019); Fang et al.
-(2020).
-While existing approaches predominantly focus on modeling user behavior through interaction histories, a critical challenge remains unaddressed: real-world interaction sequences inherently contain noise from incidental clicks and transient interests that diverge from user stable preferences.
+To address this, we propose Learned Item Shortcuts for Sequential Recommendation (LISRec), a novel framework that explicitly captures stable preferences by extracting personalized semantic shortcuts from historical interactions.
+LISRec first learns task-agnostic semantic representations to assess item similarities, then constructs a personalized semantic graph over all user-interacted items.
+By identifying the maximal semantic connectivity subset within this graph, LISRec selects the most representative items as semantic shortcuts to guide user preference modeling.
+This focused representation filters out irrelevant actions while preserving the diversity of genuine interests.
+Experimental results on the Yelp and Amazon Product datasets illustrate that LISRec achieves a 13% improvement over baseline recommendation models, showing its effectiveness in capturing stable user interests.
+Further analysis indicates that shortcut-based histories better capture user preferences, making more accurate and relevant recommendations.
+All codes and datasets are available at https://github.com/NEUIR/LISRec.
+Sequential Recommendation, Task-agnostic Item Representation, User Preference Modeling, Text-based Recommendation
+copyright: acmlicensedjournalyear: 2026doi: xxxxxxx.xxxxxxxconference: Proceedings of the 32nd SIGKDD Conference on Knowledge Discovery and Data Mining; August 9-13, 2026; Jeju, Koreaisbn: 979-x-xxxx-xxxx-x/xxxx/xxccs: Information systems Recommender systems
+Figure 1.
+Illustration of Our LISRec Method.
+LISRec constructs a personalized semantic interaction graph and extracts shortcuts, representing dominant item subsets that provide denoised user representations for preference modeling.
+1. Introduction
+Sequential recommendation systems play a pivotal role in modern digital platforms by dynamically aligning recommendations with evolving user preferences, thereby satisfying user needs and mitigating information overload (wang2019sequential; fang2020deep).
+While existing approaches predominantly focus on modeling user behavior through interaction histories, a critical challenge remains unaddressed: real-world interaction sequences inherently contain noise from incidental clicks that diverge from user stable preferences.
 As shown in Figure 1, a user consistently purchasing LEGO products might occasionally click on unrelated items like gaming consoles due to momentary curiosity.
 Such noisy interactions dilute the signal of genuine preferences, leading to suboptimal recommendation results.
-Figure 1: Illustration of Our ConsRec Method. ConsRec constructs a user-interacted item graph to denoise items for user preference modeling.
-Current methodologies primarily follow two paradigms: item-ID based modeling Kang and McAuley (2018); Sun et al.
-(2019) and item-content based modeling Hou et al.
-(2022); Liu et al.
-(2023).
+Current methodologies primarily follow two paradigms: item-ID based modeling (Wang2018self; Sun2019Bert4rec) and item-content based modeling (Hou2022uni; liu2023text).
 While these approaches learn representations from interaction sequences, they indiscriminately incorporate all items during training, implicitly assuming uniform relevance across interactions.
-This oversight amplifies noise propagation, as models struggle to distinguish between core preferences and transient outliers Lin et al.
-(2023b); Zhang et al.
-(2024a).
-Recent attempts to address noise either adjust item weights or rely on implicit contrastive learning Yang et al.
-(2023b), but these methods lack explicit mechanisms to identify and eliminate noise, limiting their interpretability.
-To model the consistent preference of users, we propose the User-Consistent Preference-based Sequential Recommendation System (ConsRec), which models user stable preferences by denoising their interacted items.
-Specifically, ConsRec constructs a user-interacted item graph by collecting all items clicked by a specific user and connecting item pairs with a higher semantic similarity.
-Then, ConsRec extracts the maximum connected subgraph from the user-interacted item graph to identify the items that represent the consistent preferences of users.
-These identified items of the maximum connected subgraph are from the densest semantic cluster, effectively helping to model stable preferences.
-Our experimental results on Amazon Product and Yelp datasets demonstrate the effectiveness of ConsRec by achieving a 13% improvement over different baseline models.
-ConsRec can also broaden its effectiveness to different recommendation architectures, showing its generalization ability.
-Further analysis shows that the retained item sets not only exhibit higher intra-cluster semantic similarity but also better match with user preferences in predicting the next item, confirming the robustness of ConsRec in broadening its effectiveness to different recommendation scenarios.
-Our work bridges the gap between semantic understanding and graph-based noise filtering, offering a principled solution to enhance sequential recommendation performance.
-2 Related Work
+This oversight amplifies noise propagation, as models struggle to distinguish between core preferences and transient outliers (lin2023self; zhang2024ssdrec).
+Recent attempts to address noise either adjust item weights (wang2021denoising), or rely on implicit contrastive learning (yang2023debiased), but these methods lack explicit mechanisms to identify and eliminate noise, limiting their interpretability.
+To address these challenges, we propose Learned Item Shortcuts for Sequential Recommendation (LISRec), a novel framework that explicitly captures stable user preferences by extracting personalized semantic shortcuts from interaction histories.
+Specifically, LISRec first learns task-agnostic semantic representations to estimate similarities between items, and then constructs a personalized semantic graph using all items previously clicked by the user, based on pairwise item similarities.
+Next, LISRec identifies a Maximal Semantic Connectivity subset of the personalized semantic graph that best captures the user’s underlying preferences, serving as semantic shortcuts to guide the user behavior modeling for recommendation.
+By focusing on these highly representative items, LISRec effectively filters out irrelevant interactions while preserving the diversity of the user’s genuine interests.
+Our experimental results on the Amazon Product and Yelp datasets demonstrate that LISRec achieves up to 13% improvement over a range of baseline models.
+Our framework generalizes well across different recommendation architectures, highlighting its flexibility.
+Further analysis shows that the extracted shortcuts consist of semantically similar items, and that these subsets are more predictive of user preferences in next-item recommendation.
+Overall, LISRec bridges the gap between semantic representation and noise filtering, providing a principled and generalizable solution for improving sequential recommendation performance.
+2. Related Work
 Sequential recommendation systems aim to model user behaviors by leveraging user-item interaction histories.
-Early studies rely on the Markov Chain assumption He and McAuley (2016) and Matrix Factorization techniques Rendle (2010) for item prediction.
-To better represent user preferences, recent works have primarily focused on ID-based recommendation modeling, which encodes the randomly initialized ID embeddings of user-interacted items using various neural architectures.
-These architectures include Recurrent Neural Networks (RNNs) Hidasi et al.
-(2016); Liu et al.
-(2016); Donkers et al.
-(2017); Yang et al.
-(2017), Convolutional Neural Networks (CNNs) Tang and Wang (2018); Yuan et al.
-(2019); Yan et al.
-(2019), Graph Neural Networks (GNNs) Chang et al.
-(2021); Zhang et al.
-(2022b); Cai et al.
-(2023), and self-attention mechanisms Kang and McAuley (2018); Sun et al.
-(2019); Du et al.
-(2023); Yang et al.
-(2023a).
-Instead of representing items using IDs, existing methods focus more on item content-based recommendation systems, leveraging Pretrained Language Models (PLMs) or Large Language Models (LLMs) to enhance item representations by learning textual semantics from their side information Ding et al.
-(2021); Yuan et al.
-(2023); Liu et al.
-(2023); Chen et al.
-(2024).
-These approaches Liu et al.
-(2023); Chen et al.
-(2024) encode the textual representations of items and further train language models using user-item interaction signals.
-TASTE Liu et al.
-(2023) verbalizes both items and user interaction histories, utilizing the T5 model Raffel et al.
-(2020) to align user and item representations.
-Leveraging the emergent abilities of LLMs Huang and Chang (2023); Lin et al.
-(2023a), HLLM Chen et al.
-(2024) introduces a hierarchical architecture for sequential recommendation comprising the item LLM and the user LLM.
-User interaction sequences often contain noise, such as accidental clicks on irrelevant items Lin et al.
-(2023b).
-Existing methods focus on self-supervised denoising, which identifies noise by learning the dependencies among items from user-item interactions Lin et al.
-(2023b); Zhang et al.
-(2024a).
-One research direction adjusts the weight distribution of items within sequences to mitigate the influence of noise, rather than directly excluding noisy items Zhou et al.
-(2022); Chen et al.
-(2022); Yang et al.
-(2023b).
-The other direction Zhang et al.
-(2023, 2022a); Sun et al.
-(2021); Qin et al.
-(2021) focuses on directly identifying and removing noisy items within interaction sequences.
-These methods predominantly focus on item ID-based denoising approaches, which require sufficient user-item interaction signals to learn effective user and item representations Lam et al.
-(2008); Schein et al.
-(2002); Pan et al.
-(2019); Zhang et al.
-(2021).
-However, they overlook the significant role of text-based semantic information in identifying consistent preferences within the user-interacted item sequence Liu et al.
-(2023); Yuan et al.
-(2023).
-Figure 2: The Model Architecture of ConsRec.
-3 Methodology
-This section introduces the User-Consistent Preference-based Sequential Recommendation System (ConsRec).
-We begin by presenting the preliminaries of the sequential recommendation model (Sec.
+Early studies rely on the Markov Chain assumption (he2016fusing) and Matrix Factorization techniques (Rendle2010Factor) for item prediction.
+To better represent user preferences, recent works have primarily focused on ID-based modeling, which encodes the randomly initialized ID embeddings of user-interacted items using various neural architectures.
+These architectures include Recurrent Neural Networks (RNNs) (Hidasi2015session; liu2016context; donkers2017sequential; yang2017neural), Convolutional Neural Networks (CNNs) (Jiaxi2018Personalized; Fajie2018asc; yan2019cosrec), Graph Neural Networks (GNNs) (Chang2021graph; zhang2022dynamic; cai2023lightgcl), and self-attention mechanisms (Wang2018self; Sun2019Bert4rec; du2023frequency; yang2023collaborative).
+Instead of representing items using IDs, existing methods focus more on item content-based recommendation systems, leveraging Pretrained Language Models (PLMs) or Large Language Models (LLMs) to enhance item representations by learning textual semantics from their side information (Ding2021zero; yuan2023go; liu2023text; chen2024hllm).
+These approaches (liu2023text; chen2024hllm) encode the textual representations of items and further train language models using user-item interaction signals.
+TASTE (liu2023text) verbalizes both items and user interaction histories, utilizing the T5 model (raffel2020exploring) to align user and item representations.
+Leveraging the emergent abilities of LLMs (huang2023towards; lin2023can), HLLM (chen2024hllm) introduces a hierarchical architecture for sequential recommendation comprising the item LLM and the user LLM.
+Figure 2. The Model Architecture of LISRec.
+User interaction sequences in recommendation systems often contain noise, such as accidental clicks or short-term behaviors that do not reflect users’ true preferences (lin2023self).
+To address this, several methods reduce the influence of noisy interactions by monitoring training dynamics (e.g., high prediction loss or unstable gradients), including adaptive sample reweighting (wang2021denoising), label correction via double correction (he2024double), and user-specific denoising strategies (zhang2025personalized).
+Beyond loss-based methods, self-supervised denoising approaches (lin2023self; zhang2024ssdrec) learn item dependencies to identify noisy signals, while other works adjust item weights based on estimated importance (zhou2022filter; chen2022denoising; yang2023debiased).
+However, these approaches typically treat denoising as an implicit process without directly removing noisy interactions.
+Another line of research (zhang2023denoising; zhang2022hierarchical; sun2021does; qin2021world) focuses on explicitly filtering out noise by identifying and removing low-quality items from interaction sequences.
+These methods predominantly rely on item ID-based representations, which require extensive user-item interaction histories to be effective (lam2008addressing; schein2002methods; pan2019warm; zhang2021language).
+Consequently, they often overlook the potential of text-based semantic information for identifying preference-consistent behaviors (liu2023text; yuan2023go).
+3. Methodology
+This section introduces LISRec, a framework for robustly modeling user preferences by extracting semantic shortcuts in a latent space.
+As shown in Figure 2, we first describe the overall architecture of LISRec, which identifies preference-consistent shortcuts to construct denoised user representations (Sec.
 3.1).
-As illustrated in Figure 2, ConsRec employs a filtering mechanism to denoise user-interacted items by identifying user-consistent preferences using item similarities (Sec.
-3.2) and then learning tailored item representations for similarity estimation using instruction tuning (Sec.
-3.3).
-3.1 Preliminary of Sequential Recommendation
+We then detail the shortcut extraction process, which identifies the Maximal Semantic Connectivity within each user’s personalized item graph (Sec.
+3.2).
+3.1. The Framework of LISRec
 The sequential recommendation task aims to predict the item $v_{t}$ that satisfies the preference of the user $\mathcal{U}$ at the $t$-th time step according to the user-item interaction history $\mathcal{H}$:
-$\mathcal{H}\rightsquigarrow M_{\text{Rec}}\rightsquigarrow v_{t},$ (1)
-where the user-item interaction history is represented as =$\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$.
+(1) $\small\mathcal{H}\rightsquigarrow M_{\text{Rec}}\rightsquigarrow v_{t},$
+where the user-item interaction history is represented as $\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$.
 The recommendation model $M_{\text{Rec}}$ estimates the relevance between the user-item interaction history $\mathcal{H}$ and an item $v$ by calculating the ranking probability $P(v|\mathcal{H})$:
-$P(v|\mathcal{H})=\text{Softmax}_{v}(\vec{\mathcal{H}}\cdot\vec{v}),$ (2)
+(2) $\small P(v|\mathcal{H})=\text{Softmax}_{v}(\vec{\mathcal{H}}\cdot\vec{v}),$
 where $\cdot$ denotes the dot product operation.
-Then we follow Liu et al.
-(2023) to get the representations $\vec{v}$ and $\vec{\mathcal{H}}$ for the candidate item $v$ and the user-item interaction history $\mathcal{H}$.
-Item Representation. To fully use the semantic information of items, we verbalize the item $v$ using its identifier and $k$ associated attributes <Attr> through the following template:
-$X(v)=\text{id:}v(\text{id})\dots\text{<Attr>}_{k}:v(\text{<Attr>}_{k}),$ (3)
+Then we follow liu2023text to get the representations $\vec{v}$ and $\vec{\mathcal{H}}$ for the candidate item $v$ and the user-item interaction history $\mathcal{H}$.
+3.1.1. Item Representation.
+To fully use the semantic information of items, we verbalize the item $v$ using its identifier and $k$ associated attributes <Attr> through the following template:
+(3) $\small X(v)=\texttt{id:}v(\texttt{id})\dots\texttt{<Attr>}_{k}:v(\texttt{<Attr>}_{k}),$
+where $\texttt{<Attr>}_{k}$ is the name of the $k$-th attribute, while $v(\texttt{id})$ and $v(\texttt{<Attr>}_{i})$ correspond to the text descriptions of the item identifier and the $i$-th attribute, respectively.
+Then we encode each item $v$ into an embedding $\vec{v}$ using the T5 model (raffel2020exploring):
+(4) $\small\vec{v}=\text{T5}(X(v)).$
+3.1.2. Modeling User Preferences via Shortcut Extraction
+Users’ interaction histories ($\mathcal{H}$) often contain a mixture of long-term interests, short-term preferences, and incidental click behaviors.
+To model user intent using truly essential items, LISRec tries to model preference-consistent behaviors by building semantic shortcuts in the user-interacted item trajectory to omit some noisy items.
+To identify these shortcuts, we introduce a semantic shortcut extractor $M_{\text{Filter}}$, which selectively retrieves a subset of dominant interactions from the full history $\mathcal{H}$.
+These shortcuts, denoted as $\mathcal{H}_{\text{Shortcut}}$, form a condensed yet highly informative sketch of user preferences in the latent semantic space.
+This distilled representation is then passed to the recommendation module $M_{\text{Rec}}$ for more effective user behavior modeling.
+The user representation can be formalized as:
+(5) $\small\mathcal{H}\xrightarrow{M_{\text{Filter}}}\mathcal{H}_{\text{Shortcut}},$
+where $\mathcal{H}_{\text{Shortcut}}\subseteq\mathcal{H}$ denotes the extracted shortcut subset, containing interactions that are semantically aligned with the user’s underlying intent.
+The semantic shortcut extractor $M_{\text{Filter}}$ are detailed in Section 3.2.
+To construct the user representation, we first verbalize the extracted shortcut set $\mathcal{H}_{\text{Shortcut}}$ by concatenating the textual representations (Eq. 3) of the selected items:
+(6) $\small X(\mathcal{H}_{\text{Shortcut}})=\mathop{\|}_{v_{i}\in\mathcal{H}_{\text{Shortcut}}}X(v_{i}),$
+where $\|$ denotes the concatenation operation over the subset of items in $\mathcal{H}_{\text{Shortcut}}$, which may be non-contiguous and unordered with respect to the original history.
+The concatenated text is then encoded using the T5 model to obtain a semantic representation of the user’s preference:
+(7) $\small\vec{\mathcal{H}}_{\text{Shortcut}}=\text{T5}(X(\mathcal{H}_{\text{Shortcut}})).$
+3.2. Extracting Preference Shortcuts via Maximal Semantic Connectivity
+To robustly model stable user preferences from noisy interaction histories $\mathcal{H}$, we introduce a semantic shortcut extractor $M_{\text{Filter}}$, which identifies the most representative subset of interactions for preference modeling.
+Specifically, LISRec first encodes items into a task-agnostic semantic space to estimate pairwise item similarity, then constructs a personalized semantic graph for each user and extracts the maximally connected component as the preference-consistent shortcut.
+This process effectively filters out noisy or inconsistent behaviors during user representation construction.
+3.2.1. Task-Agnostic Semantic Item Encoding
+To obtain semantically meaningful item representations suitable for preference-level reasoning, we train a language model on item textual descriptions using the instruction tuning based item representation method.
+For item semantic representation, each item $v$ is represented purely through its descriptive attributes—excluding any item ID—to improve generalization and reduce dependency on sparse or domain-specific identifiers:
+(8) $\small\tilde{X}(v)^{\prime}=\texttt{<Attr>}_{1}:v(\texttt{<Attr>}_{1})\ \dots\ \texttt{<Attr>}_{k}:v(\texttt{<Attr>}_{k}).$
+Building upon the Next Item Prediction (NIP) objective (Sun2019Bert4rec), we serialize each user history $\mathcal{H}=\{v_{1},\dots,v_{t-1}\}$ into a concatenated text sequence:
+(9) $\small\tilde{X}^{\prime}(\mathcal{H})=\tilde{X}(v_{t-1})^{\prime};\dots;\tilde{X}(v_{1})^{\prime}.$
 […]
-$\vec{v}=\text{T5}(X(v)).$ (4)
-User Representation.
-To represent the user preference, we verbalize the user-item interaction history $\mathcal{H}$ as a concatenation of the text representations (Eq.
-3) of the items in $\mathcal{H}$:
-$X(\mathcal{H})=X(v_{t-1});\dots;X(v_{1}),$ (5)
-where ; denotes the concatenation operation. Then we encode $\mathcal{H}$ into embeddings $\vec{\mathcal{H}}$ using the T5 model, which shares the same parameters with the encoder model in Eq. 4:
-$\vec{\mathcal{H}}=\text{T5}(X(\mathcal{H})).$ (6)
-3.2 Denoising User-Interacted Items Using User-Consistent Preference
-In user-item interactions, users may occasionally display sudden behavioral changes or lose interest in previously preferred items.
-Such fluctuations introduce noise into the interaction history, complicating accurate user preference modeling.
-To tackle this issue, we incorporate a denoising module ($M_{\text{Filter}}$) to identify and retain items that reflect consistent user preferences.
-Then we train the sequential recommendation module ($M_{\text{Rec}}$) by constructing datasets using the denoised user-item interaction history $\mathcal{H}_{\text{Filter}}$ and the $t$-th step item $v_{t}$.
-After training, the recommendation module is used to predict the item $v_{t}$ based on $\mathcal{H}_{\text{Filter}}$:
-$\mathcal{H}\rightsquigarrow M_{\text{Filter}}\xrightarrow{\mathcal{H}_{\text{Filter}}}M_{\text{Rec}}\rightsquigarrow v_{t},$ (7)
-where $\mathcal{H}_{\text{Filter}}\subseteq\mathcal{H}$ represents the subset of items retained to capture consistent preferences of the user.
-To implement the noise filtering model $M_{\text{Filter}}$, we construct a user-interacted item graph by using the item similarity and then extract the maximum connected component from the graph to obtain the set of denoise items $\mathcal{H}_{\text{Filter}}$.
-User-Interacted Item Graph. We first introduce the user-interacted item graph, which is used to distinguish some outlier items that are different from the user’s general preference.
-$G_{\mathcal{U}}=(\mathcal{V},\mathcal{E}),$ (8)
-where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ represents all items that have been interacted by user $\mathcal{U}$. The $\mathcal{E}$ is an edge matrix:
-=$\mathcal{E}=\{e_{ij},1\leq i\leq n,1\leq j\leq n\},$ (9)
-where the element $e_{ij}$ indicates whether the $i$-th node $v_{i}$ and the $j$-th node $v_{j}$ are connected with an edge. The edge is established according to the similarity score $S_{ij}$:
-$e_{ij}=\begin{cases}0&\text{if }S_{ij}<\tau,\\ 1&\text{if }S_{ij}\geq\tau,\end{cases}$ (10)
-where $\tau$ is the threshold of the similarity.
-$e_{ij}=0$ indicates that the $i$-th node $v_{i}$ and the $j$-th node $v_{j}$ are not connected, while $e_{ij}=1$ indicates that the nodes are connected.
-To estimate the similarity between $v_{i}$ and $v_{j}$, we encode all items in the graph node set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ into embeddings $\{\vec{v}_{1},\dots,\vec{v}_{n}\}$.
-Subsequently, we calculate the similarity score $S_{ij}$ using cosine similarity:
-$S_{ij}=\cos(\vec{v}_{i},\vec{v}_{j}).$ (11)
-Then we will introduce how to get the item embedding for similarity calculation in Sec. 3.3.
-Denoising User-Interacted Items Using the Maximum Connected Component.
-To emphasize stronger connections among items, we aim to identify the connected components of the user-interacted item graph $G_{\mathcal{U}}$.
-Specifically, we perform a Breadth-First Search (BFS) to find a set of disjoint connected components:
-$\mathcal{C}=\{C_{1},\dots,C_{l}\},$ (12)
-where $l$ is the number of connected components, and each $C_{i}$ is a subgraph of $G_{\mathcal{U}}$.
-For each connected component $C_{i}$, we compute its size, defined as the number of nodes $|\mathcal{V}(C_{i})|$.
-The maximum connected subgraph $C_{\text{max}}$ is then selected as:
-$C_{\text{max}}=\arg\max_{C_{i}}|\mathcal{V}(C_{i})|.$ (13)
-These user-interacted items that belong to the node set $\mathcal{V}(C_{\text{max}})$ of the maximum connected subgraph $C_{\text{max}}$ are retained to form the denoised interaction sequence $\mathcal{H}_{\text{Filter}}$:
-$\mathcal{H}_{\text{Filter}}=\{v_{j}\mid v_{j}\in\mathcal{H}\land v_{j}\in\mathcal{V}(C_{\text{max}})\}.$ (14)
-This retained interaction sequence $\mathcal{H}_{\text{Filter}}$ is then passed to the recommendation module $M_{\text{Rec}}$ to predict the item at the $t$-th step.
-More detailed theoretical explanations are presented in Appendix A.7.
-3.3 Universal Item Representation Learning for User-Interacted Graph Construction
-To obtain item representations that are effective for constructing edges in the graph $G_{u}$, we adopt instruction tuning Chung et al.
-(2024); Wang et al.
-(2023) to train language models such as T5 Raffel et al.
-(2020) based on their textual inputs.
-This approach differs fundamentally from existing denoising-based methods Yuan et al.
-(2021a); Zhang et al.
-(2022a); Yang et al.
-(2023b); Zhang et al.
-(2024a), as it leverages supervised signals from different downstream recommendation tasks.
-To learn text-based item representations for various recommendation tasks $\mathcal{T}$, we represent each item $v$ only through its textual semantic information.
-Unlike prior works relying on ID-based representations (Eq.
-3), we verbalize the item only using its descriptive attributes and exclude the item ID, enhancing the generalization ability of the item representation model:
-$\tilde{X}(v)^{\prime}=\text{<Attr>}_{1}:v(\text{<Attr>}_{1})\dots\text{<Attr>}_{k}:v(\text{<Attr>}_{k}).$ (15)
-To fully use the user-item interaction =${\mathcal{H}}=\{v_{1},v_{2},\dots,v_{t-1}\}$, we follow previous work in recommendation model pretraining Sun et al.
-(2019) and use the Next Item Prediction (NIP) task to train item representations (Eq.
-2).
-Specifically, we verbalize the user-item interaction history $\mathcal{H}$ as a concatenation of verbalized items =$\mathcal{H}=\{v_{1},v_{2},\dots,v_{t-1}\}$:
-$\tilde{X}^{\prime}(\mathcal{H})=\tilde{X}(v_{t-1})^{\prime};\dots;\tilde{X}(v_{1})^{\prime},$ (16)
-Then we use the instruction to illustrate the Next Item Prediction task: “This is $\mathcal{T}$ dataset.
-Here is the visit history list of the user: $\tilde{X}(\mathcal{H})^{\prime}$, recommend next item”, aiming to predict the $t$-th step item $v_{t}$.
-4 Experimental Methodology
+(10) $\small S_{ij}=\cos(\vec{v}_{i},\vec{v}_{j}).$
+This representation approach is fully text-driven, enabling the item representation model to operate across item vocabularies and different tasks while capturing semantic regularities grounded in language.
+3.2.2. Modeling Stable Preferences with Personal Item Graphs
+To extract the user’s core preference signals from noisy interaction histories, we construct a personalized semantic interaction graph for each user $\mathcal{U}$, where edges capture semantic coherence between interacted items:
+(11) $\small G_{\mathcal{U}}=(\mathcal{V},\mathcal{E}),$
+where $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ includes all items interacted by user $\mathcal{U}$ in all previous sessions, and $\mathcal{E}$ denotes the edge set encoding semantic connections.
+An undirected edge $e_{ij}\in\mathcal{E}$ is created between items $v_{i}$ and $v_{j}$ if their embedding similarity exceeds a threshold $\tau$:
+(12) $\small e_{ij}=\begin{cases}0&\text{if }S_{ij}<\tau,\\ 1&\text{if }S_{ij}\geq\tau,\end{cases}$
+where $e_{ij}=0$ is the cosine similarity between the embeddings of $v_{i}$ and $v_{j}$ are not connected (Eq.
+10).
+This design naturally filters out semantically dissimilar and outlier items with respect to user preferences.
+To identify the user’s dominant preference signal, we extract the largest connected component from the graph $G_{\mathcal{U}}$:
+(13) $\small C_{\text{max}}=\arg\max_{C_{i}}|\mathcal{V}(C_{i})|,$
+and define the corresponding item subset as the preference shortcut:
+(14) $\small\mathcal{H}_{\text{Shortcut}}=\{v_{j}\mid v_{j}\in\mathcal{H}\land v_{j}\in\mathcal{V}(C_{\text{max}})\}.$
+This shortcut subset serves as a semantically consistent and compact proxy for the user’s intent.
+By selectively modeling only the maximally connected region in the semantic graph, we avoid overfitting to noisy interactions while mitigating over, ensuring richer and more robust preference representations for downstream recommendation.
+4. Experimental Methodology
 This section describes the datasets, evaluation metrics, baseline models, and implementation details used in our experiments.
 Datasets.
-We utilize four sequential recommendation datasets–Beauty, Sports, Toys McAuley et al.
-(2015), and Yelp–to train and evaluate both the sequential recommendation module and the item filtering module.
+We utilize four sequential recommendation datasets–Beauty, Sports, and Toys from Amazon Product (mcauley2015image), and Yelp–to train and evaluate both the sequential recommendation module and the semantic shortcut extractor.
 More details are shown in Appendix A.2.
 Evaluation Metrics.
-We adopt Recall@10/20 and NDCG@10/20 to evaluate the recommendation performance, which is consistent with prior works Xie et al.
-(2022); Liu et al.
-(2023); Chen et al.
-(2024).
+We adopt Recall@10/20 and NDCG@10/20 to evaluate the recommendation performance, which is consistent with prior works (Xie2022DIF-SR; liu2023text; chen2024hllm).
 Statistical significance is determined using a permutation test with $P<0.05$.
-Baselines. We compare ConsRec with several widely used methods for sequential recommendation, including vanilla sequential recommendation models and some item denoising techniques.
-Vanilla Sequential Recommendation Models. We compare ConsRec with both ID-based and content-based recommendation models.
-First, ConsRec is evaluated against seven ID-based recommendation models.
-GRU4Rec Hidasi et al.
-(2016) employs Recurrent Neural Networks (RNNs) to model user-item interaction sequences, while SASRec Kang and McAuley (2018) and Bert4Rec Sun et al.
-(2019) leverage self-attention to capture user preferences from sequential interactions.
+Baselines. We compare LISRec with several widely used methods for sequential recommendation, including vanilla sequential recommendation models and some item denoising techniques.
+Vanilla Sequential Recommendation Models.
+We compare LISRec with both ID-based and content-based recommendation models.
+First, LISRec is evaluated against seven ID-based recommendation models.
+GRU4Rec (Hidasi2015session) employs Recurrent Neural Networks (RNNs) to model user-item interaction sequences, while SASRec (Wang2018self) and Bert4Rec (Sun2019Bert4rec) leverage self-attention to capture user preferences from sequential interactions.
 Additionally, we examine models that incorporate item-side information.
-S3Rec Zhou et al.
-(2020) enhances self-attention modules through four self-supervised pretraining strategies to better model relationships among attributes, items, and users.
-ICAI-SR Yuan et al.
-(2021b) represents item-attribute relationships using a heterogeneous graph to improve relevance modeling, whereas NOVA Liu et al.
-(2021a) integrates item attributes as side information via attention modules.
-DIF-SR Xie et al.
-(2022) introduces a non-invasive attention mechanism to fuse item attribute information into user behavior modeling effectively.
-Furthermore, ConsRec is compared with four content-based recommendation models.
-CoWPiRec Yang et al.
-(2023a) optimizes Pretrained Language Models (PLMs) to enhance item representation by aligning it with collaborative filtering representations derived from a word graph.
-TASTE Liu et al.
-(2023) leverages T5 Raffel et al.
-(2020) to directly encode the text representations of both users and items for matching items with user intentions.
-DWSRec Zhang et al.
-(2024b) leverages both fully whitened and relaxed whitened item representations as dual views for effective recommendations.
-Lastly, HLLM Chen et al.
-(2024) employs Large Language Models (LLMs) to generate item representations, subsequently utilizing a hierarchical architecture to predict the next item.
+S3Rec (Zhou2020s3) enhances self-attention modules through four self-supervised pretraining strategies to better model relationships among attributes, items, and users.
+ICAI-SR (yuan2021icai) represents item-attribute relationships using a heterogeneous graph to improve relevance modeling, whereas NOVA (Liu2021nova) integrates item attributes as side information via attention modules.
+DIF-SR (Xie2022DIF-SR) introduces a non-invasive attention mechanism to fuse item attribute information into user behavior modeling effectively.
+Furthermore, LISRec is compared with four content-based recommendation models.
+CoWPiRec (yang2023collaborative) optimizes Pretrained Language Models (PLMs) to enhance item representation by aligning it with collaborative filtering representations derived from a word graph.
+TASTE (liu2023text) leverages T5 (raffel2020exploring) to directly encode the text representations of both users and items for matching items with user intentions.
+DWSRec (zhang2024dual) leverages both fully whitened and relaxed whitened item representations as dual views for effective recommendations.
+Lastly, HLLM (chen2024hllm) employs Large Language Models (LLMs) to generate item representations, subsequently utilizing a hierarchical architecture to predict the next item.
+Table 1.
+Overall Performance of LISRec.
+The best results are marked in bold, while the second-best results are underlined.
+The HLLM model is reproduced using a 1B-scale LLM.
+$\dagger$} and $\ddagger$} denote statistically significant improvements over $\text{DIF-SR}^{\dagger}$ and $\text{TASTE}^{\ddagger}$, respectively.
+Dataset Metrics Item ID based Recommendation Models Text based Recommendation Models
+GRU4Rec Bert4Rec SASRec S3Rec NOVA ICAI-SR DIF-SR CoWPiRec DWSRec HLLM TASTE LISRec
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Item Denoising for Sequential Recommendation.
-We also evaluate ConsRec against three baselines that either explicitly remove noisy items or implicitly address item noise through denoising modeling.
-All these models are implemented based on the item-ID base recommendation model, such as Bert4Rec Sun et al.
-(2019).
-For methods that explicitly remove noisy items, DSAN Yuan et al.
-(2021a) learns a target item embedding and applies an adaptively sparse transformation function to filter out noisy interactions.
-Similarly, HSD Zhang et al.
-(2022a) identifies and excludes noisy interactions by detecting inconsistencies with user interests, leveraging the sequential order of recorded interactions.
-In contrast, DCRec Yang et al.
-(2023b) employs a debiased contrastive learning framework to implicitly mitigate the impact of noise.
-Dataset Metrics Item ID based Recommendation Models Text based Recommendation Models
-GRU4Rec Bert4Rec SASRec S3Rec NOVA ICAI-SR DIF-SR CoWPiRec DWSRec HLLM TASTE ConsRec
-Beauty R@10 0.0530 0.0529 0.0828 0.0868 0.0887 0.0879 0.0908 0.0837 0.0989 0.0887 0.1030 $\text{}^{\dagger\ddagger}$
-R@20 0.0839 0.0815 0.1197 0.1236 0.1237 0.1231 0.1284 0.1280 0.1496 0.1238 0.1550 $\text{}^{\dagger\ddagger}$
-N@10 0.0266 0.0237 0.0371 0.0439 0.0439 0.0439 0.0446 0.0405 0.0470 0.0499 0.0517 $\text{}^{\dagger\ddagger}$
-N@20 0.0344 0.0309 0.0464 0.0531 0.0527 0.0528 0.0541 0.0517 0.0597 0.0588 0.0649 $\text{}^{\dagger\ddagger}$
-Sports R@10 0.0312 0.0295 0.0526 0.0517 0.0534 0.0527 0.0556 0.0588 0.0629 0.0555 0.0633 $\text{}^{\dagger\ddagger}$
-R@20 0.0482 0.0465 0.0773 0.0758 0.0759 0.0762 0.0800 0.0892 0.0958 0.0816 0.0964 $\text{}^{\dagger\ddagger}$
-N@10 0.0157 0.0130 0.0233 0.0249 0.0250 0.0243 0.0264 0.0300 0.0313 0.0295 0.0338 $\text{}^{\dagger\ddagger}$
-N@20 0.0200 0.0173 0.0295 0.0310 0.0307 0.0302 0.0325 0.0376 0.0396 0.0360 0.0421 $\text{}^{\dagger\ddagger}$
-Toys R@10 0.0370 0.0533 0.0831 0.0967 0.0978 0.0972 0.1013 0.0502 0.0967 0.0928 0.1232 $\text{}^{\dagger\ddagger}$
-R@20 0.0588 0.0787 0.1168 0.1349 0.1322 0.1303 0.1382 0.0739 0.1472 0.1306 0.1789 $\text{}^{\dagger\ddagger}$
-N@10 0.0184 0.0234 0.0375 0.0475 0.0480 0.0478 0.0504 0.0272 0.0457 0.0520 0.0640 $\text{}^{\dagger}$
-N@20 0.0239 0.0297 0.0460 0.0571 0.0567 0.0561 0.0597 0.0331 0.0585 0.0616 0.0780 $\text{}^{\dagger}$
-Yelp R@10 0.0361 0.0524 0.0650 0.0589 0.0681 0.0663 0.0698 0.0657 0.0702 0.0659 0.0738 $\text{}^{\dagger\ddagger}$
-R@20 0.0592 0.0756 0.0928 0.0902 0.0964 0.0940 0.1003 0.0959 0.1103 0.0955 0.1156 $\text{}^{\dagger\ddagger}$
-N@10 0.0184 0.0327 0.0401 0.0338 0.0412 0.0400 0.0419 0.0416 0.0382 0.0407 0.0397 $\text{}^{\ddagger}$
-N@20 0.0243 0.0385 0.0471 0.0416 0.0483 0.0470 0.0496 0.0497 0.0482 0.0504 0.0502 $\text{}^{\dagger\ddagger}$
-[…]
+We also evaluate LISRec against three baselines that either explicitly remove noisy items or implicitly address item noise through denoising modeling.
+All these models are implemented based on the item-ID base recommendation model, such as Bert4Rec (Sun2019Bert4rec).
+For methods that explicitly remove noisy items, DSAN (yuan2021dual) learns a target item embedding and applies an adaptively sparse transformation function to filter out noisy interactions.
+Similarly, HSD (zhang2022hierarchical) identifies and excludes noisy interactions by detecting inconsistencies with user interests, leveraging the sequential order of recorded interactions.
+In contrast, DCRec (yang2023debiased) employs a debiased contrastive learning framework to implicitly mitigate the impact of noise.
 Implementation Details.
-Our ConsRec is implemented using OpenMatch Liu et al.
-(2021b); Yu et al.
-(2023) and optimized using in-batch negatives.
-Both the item representation model in the filter module and the recommendation module are initialized from the T5-base checkpoint [^1] https://huggingface.co/google-t5/t5-base in Huggingface Transformers Wolf et al.
-(2019).
-The item representation model in the filter module is trained using the Adam optimizer with a learning rate of 1e-4 and a batch size of 32.
-For the recommendation module, we follow the settings in Liu et al.
-(2023), which is trained using the Adam optimizer with a learning rate of 1e-4, a warm-up proportion of 0.1, and a batch size of 8.
-We set the filtering threshold $\tau$ to 0.7, based on the hyperparameter sensitivity analysis reported in Appendix A.4.
+LISRec is implemented using OpenMatch [^1] https://github.com/OpenMatch/OpenMatch (liu2021openmatch; yu2023openmatch) and optimized using in-batch negatives.
+Both the item representation model in the semantic shortcut extractor and the recommendation module are initialized from the T5-base checkpoint [^2] https://huggingface.co/google-t5/t5-base in Huggingface Transformers (wolf2019huggingface).
+The item representation model in the semantic shortcut extractor is trained using the Adam optimizer with a learning rate of 1e-4 and a batch size of 32.
+For the recommendation module, we follow the settings in liu2023text, which is trained using the Adam optimizer with a learning rate of 1e-4, a warm-up proportion of 0.1, and a batch size of 8.
+We set the filtering threshold $\tau$ to 0.7, based on the hyperparameter sensitivity analysis reported in Appendix 5.3.
 During inference, we excluded the last interaction from the user-interaction sequence when constructing the consistency graph, thereby avoiding potential data leakage.
-5 Experimental Results
-In this section, we begin by evaluating the recommendation performance of ConsRec and perform ablation studies to assess the effectiveness of denoising modules and various training strategies.
-Then, we analyze text similarities among retained items to evaluate the effectiveness of the item representation learning method.
-Finally, case studies are shown in Appendix A.6.
-5.1 Overall Performance
-In this section, we first implement ConsRec based on TASTE Liu et al.
-(2023) to evaluate its recommendation performance.
-Subsequently, we implement ConsRec with Bert4Rec Sun et al.
-(2019) to evaluate its denoising capability by comparing it against various baselines.
-Recommendation Performance.
-As shown in Table 1, we present the overall performance of ConsRec and various recommendation baseline models.
-Overall, ConsRec demonstrates its effectiveness by significantly outperforming all baselines across four datasets, leveraging TASTE Liu et al.
-(2023) as its backbone architecture.
+Additionally, we reproduce all item-ID based models and DWSRec (zhang2024dual) using implementations provided by RecBole [^3] https://recbole.io/ (zhao2021recbole), maintaining consistent settings across all models for fair comparison.
+For CoWPiRec and TASTE, we adhere to the default experimental settings as specified in their original papers.
+For HLLM, due to computational constraints, we ultilize TinyLlama-1.1B checkpoint [^4] https://huggingface.co/TinyLlama/TinyLlama-1.1B-intermediate-step-1431k-3T (zhang2024tinyllama) as the item-LLM component, and the user-LLM is implemented as a custom Llama model (touvron2023llama) with two hidden layers, trained from scratch using the following settings: learning rate of 1e-4, batch size of 2, training for 5 epochs, and a negative sampling ratio of 1024 negative samples per positive sample.
+5. Experimental Results
+In this section, we begin by evaluating the recommendation performance of LISRec and perform ablation studies to assess the effectiveness of the semantic shortcut extractor and various training strategies.
+Then, we analyze textual similarities among retained items to evaluate the effectiveness of the item representation learning method.
+Finally, several case studies are shown.
+5.1. Overall Performance
+In this section, we first implement LISRec based on TASTE (liu2023text) to evaluate its recommendation performance.
+Subsequently, we implement LISRec with Bert4Rec (Sun2019Bert4rec) to evaluate its denoising capability by comparing it against various baselines.
+5.1.1. Recommendation Performance
+As shown in Table 1, we present the overall performance of LISRec and various recommendation baseline models.
+Overall, LISRec demonstrates its effectiveness by significantly outperforming all baselines across four datasets, leveraging TASTE (liu2023text) as its backbone architecture.
 Among all baseline models, text-based recommendation models generally exhibit significantly stronger performance compared to item ID-based models.
 This superiority derives from the ability of PLMs and LLMs to effectively capture matching signals through text representations of users and items.
-Based on the TASTE backbone, ConsRec integrates a denoising module to remove noisy items, utilizing these retained user-item interactions to train the recommendation module.
+Based on the TASTE backbone, LISRec integrates a semantic shortcut extractor to remove noisy items, utilizing these retained user-item interactions to train the recommendation module.
 The evaluation results reveal a significant improvement in the recommendation performance, highlighting its effectiveness in enabling the recommendation module to better model user behavior.
-Among various recommendation tasks, ConsRec demonstrates much better performance on the Beauty and Sports datasets, while showing relatively weaker improvements on the Yelp dataset.
+Among various recommendation tasks, LISRec demonstrates much better performance on the Beauty and Sports datasets, while showing relatively weaker improvements on the Yelp dataset.
 This discrepancy may be attributed to the tendency of users to make more casual clicks driven by instant interest in these shopping scenarios.
+Table 2. Denoising Performance of LISRec. To ensure a fair comparison, all denoising methods are implemented using Bert4Rec (Sun2019Bert4rec) as the backbone model.
 Dataset Method R@10 R@20 N@10 N@20
@@
 DCRec 0.0609 0.0869 0.0317 0.0387
-ConsRec 0.0627 0.0904 0.0336 0.0406
+LISRec 0.0627 0.0904 0.0336 0.0406
 Sports DSAN 0.0107 0.0159 0.0051 0.0065
@@
 DCRec 0.0262 0.0407 0.0139 0.0175
-ConsRec 0.0275 0.0428 0.0147 0.0185
+LISRec 0.0275 0.0428 0.0147 0.0185
 Toys DSAN 0.0172 0.0258 0.0086 0.0104
@@
 DCRec 0.0443 0.0660 0.0243 0.0296
-ConsRec 0.0468 0.0691 0.0258 0.0314
+LISRec 0.0468 0.0691 0.0258 0.0314
 Yelp DSAN 0.0153 0.0218 0.0084 0.0116
@@
 DCRec 0.0387 0.0562 0.0204 0.0262
-ConsRec 0.0398 0.0576 0.0211 0.0274
-Table 2: Denoising Performance of Different Methods. To ensure a fair comparison, all denoising methods are implemented using Bert4Rec Sun et al. (2019).
-Denoising Effectiveness.
-We compare ConsRec with three state-of-the-art denoising methods–DSAN Yuan et al.
-(2021a), HSD Zhang et al.
-(2022a), and DCRec Yang et al.
-(2023b)–to evaluate its denoising effectiveness.
+LISRec 0.0398 0.0576 0.0211 0.0274
+5.1.2. Denoising Effectiveness
+We compare LISRec with three state-of-the-art denoising methods–DSAN (yuan2021dual), HSD (zhang2022hierarchical), and DCRec (yang2023debiased)–to evaluate its denoising effectiveness.
 The baselines are categorized into two groups–explicit denoising methods (DSAN and HSD) and implicit denoising methods (DCRec).
-All methods are implemented using the item ID-based recommendation model, Bert4Rec Sun et al.
-(2019).
+All methods are implemented using the item ID-based recommendation model, Bert4Rec (Sun2019Bert4rec).
 The evaluation results are shown in Table 2.
@@
 However, this implicit denoising approach necessitates model-specific training, restricting our experiments to the Bert4Rec model used in constructing the DCRec framework.
-Benefiting from explicit denoising methods, ConsRec can be applied to the item ID-based recommendation model and achieves a 4.7% improvement over all denoising baseline models, highlighting its effectiveness in filtering noisy items.
-This consistent improvement with the item ID-based model underscores the potential of extending the filtering approach of ConsRec to different architectures of recommendation models.
+Benefiting from explicit denoising methods, LISRec can be applied to the item ID-based recommendation model and achieves a 4.7% improvement over all denoising baseline models, highlighting its effectiveness in filtering noisy items.
+This consistent improvement with the item ID-based model underscores the potential of extending the filtering approach of LISRec to different architectures of recommendation models.
+Table 3.
+Ablation Study of LISRec.
+$\dagger$} and $\ddagger$} denote statistically significant improvements over $\text{LISRec{} (w/o $M_{\text{Filter}}$)}^{\dagger}$ and $\text{LISRec{} (Zero-Shot)}^{\ddagger}$.
 Dataset Method R@20 N@20
-Beauty ConsRec (Zero-Shot) 0.1631 0.0697
+Beauty LISRec (Zero-Shot) 0.1631 0.0697
 w/o $M_{\text{Filter}}$ 0.1550 0.0649
 […]
-w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
-w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
-Sports ConsRec (Zero-Shot) 0.1014 0.0420
+[…]
+[…]
+Sports LISRec (Zero-Shot) 0.1014 0.0420
 w/o $M_{\text{Filter}}$ 0.0964 0.0421
 […]
-w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
-w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
-Toys ConsRec (Zero-Shot) 0.1797 0.0786
+[…]
+[…]
+Toys LISRec (Zero-Shot) 0.1797 0.0786
 w/o $M_{\text{Filter}}$ 0.1789 0.0780
 […]
-w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
-w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
-Yelp ConsRec (Zero-Shot) 0.1158 0.0504
+[…]
+[…]
+Yelp LISRec (Zero-Shot) 0.1158 0.0504
 w/o $M_{\text{Filter}}$ 0.1156 0.0502
@@
 w/o $M_{\text{Rec}}$ Training 0.1165 0.0504
-w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
 […]
-5.2 Ablation Study
-This section presents an ablation study to evaluate the contribution of the denoising module ($M_{\text{Filter}}$) and the effectiveness of our proposed training strategy in ConsRec.
-More experimental results are shown in Appendix A.5.
+Table 4. Ablation Study on Different Filter Models in LISRec.
+Dataset Method R@20 N@20
+Beauty LISRec 0.1774 0.0772
+w/ Vanilla T5 $M_{\text{Filter}}$ 0.1761 0.0762
+w/ TASTE $M_{\text{Filter}}$ 0.1767 0.0767
+Sports LISRec 0.1082 0.0496
+w/ Vanilla T5 $M_{\text{Filter}}$ 0.1073 0.0457
+w/ TASTE $M_{\text{Filter}}$ 0.1077 0.0471
+Toys LISRec 0.1896 0.0826
+w/ Vanilla T5 $M_{\text{Filter}}$ 0.1879 0.0802
+w/ TASTE $M_{\text{Filter}}$ 0.1883 0.0811
+Yelp LISRec 0.1173 0.0509
+w/ Vanilla T5 $M_{\text{Filter}}$ 0.1169 0.0507
+w/ TASTE $M_{\text{Filter}}$ 0.1170 0.0508
+5.2. Ablation Study
+This section presents an ablation study to evaluate the contribution of the semantic shortcut extractor ($M_{\text{Filter}}$) and the effectiveness of our proposed training strategy in LISRec.
 Table 3 shows the performance of several ablated versions of our model across four datasets.
-Comparing ConsRec (w/o $M_{\text{Filter}}$) with ConsRec (Zero-Shot), we observe a consistent improvement in recommendation performance across all datasets when a zero-shot $M_{\text{Filter}}$ is applied during inference.
-This highlights the inherent benefit of even an untrained denoising module in refining user-interacted items for recommendation.
-The impact of the denoising module becomes more evident when comparing ConsRec (Zero-Shot) with ConsRec (w/o $M_{\text{Rec}}$ Training).
-Training the $M_{\text{Filter}}$ module using our instruction-based method leads to a significant improvement, demonstrating the effectiveness of learning tailored item representations through denoising.
-Specifically, ConsRec (w/o $M_{\text{Rec}}$ Training) consistently outperforms ConsRec (w/o $M_{\text{Filter}}$), indicating that a well-trained filter is crucial for recommendation systems.
-Furthermore, the full ConsRec model, which trains the recommendation module $M_{\text{Rec}}$ on the retained interactions, achieves the best performance across all datasets, with an average improvement of over 12% compared to ConsRec (Zero-Shot).
+Comparing LISRec (w/o $M_{\text{Filter}}$) with LISRec (Zero-Shot), we observe a consistent improvement in recommendation performance across all datasets when a zero-shot $M_{\text{Filter}}$ is applied during inference.
+This highlights the inherent benefit of even an untrained semantic shortcut extractor in refining user-interacted items for recommendation.
+The impact of the semantic shortcut extractor becomes more evident when comparing LISRec (Zero-Shot) with LISRec (w/o $M_{\text{Rec}}$ Training).
+Training the $M_{\text{Filter}}$ module with our instruction-based approach results in significant performance gains, demonstrating the value of learning tailored item representations for extracting semantic shortcuts.
+In particular, LISRec (w/o $M_{\text{Rec}}$ Training) consistently outperforms LISRec (w/o $M_{\text{Filter}}$), highlighting that an effective semantic shortcut extractor is essential for improved recommendation accuracy.
+Furthermore, the full LISRec model, which trains the recommendation module $M_{\text{Rec}}$ on the retained interactions, achieves the best performance across all datasets, with an average improvement of over 12% compared to LISRec (Zero-Shot).
 This substantial gain underscores the importance of aligning the recommendation module with the retained data, enabling it to better capture user preferences and behaviors from the refined interaction history.
-The consistent statistically significant improvements of ConsRec over both ConsRec (w/o $M_{\text{Filter}}$) and ConsRec (Zero-Shot) further validate the effectiveness of our complete model.
+The consistent statistically significant improvements of LISRec over both LISRec (w/o $M_{\text{Filter}}$) and LISRec (Zero-Shot) further validate the effectiveness of our complete model.
+To evaluate the effectiveness of the $M_{\text{Filter}}$ module in LISRec, we compare it with two alternatives: a vanilla T5 $M_{\text{Filter}}$ initialized from T5-base without task-specific training, and a TASTE $M_{\text{Filter}}$ (trained on recommendation data).
+As shown in Table 4, LISRec achieves the best performance across all datasets, consistently surpassing both baselines.
+The vanilla T5 $M_{\text{Filter}}$ lags behind due to its lack of adaptation to recommendation data, while the TASTE-based variant, though improved, is still limited by its exposure to only a single dataset during training.
+In contrast, our approach benefits from balanced pre-training over multiple datasets, leading to more robust and transferable item representations.
+These results highlight the importance of both domain-specific and multi-dataset training for semantic shortcut extraction in recommendation.
+(a) Bert4Rec
+(b) TASTE
+Figure 3.
+Hyperparameter Sensitivity Analysis.
+We conduct experiments on the threshold hyperparameter, $\tau$, with results derived from the Amazon Beauty dataset using the Bert4Rec 3(a) and TASTE 3(b) backbone models.
+5.3. Hyperparameter Sensitivity Analysis
+We analyze the impact of the similarity threshold $\tau$, which is a key hyperparameter in the construction of the personal item graph described in Sec.3.2.2, by varying $\tau$ from 0.2 to 0.8 while keeping all other experimental settings fixed.
+Figure 3 shows the performance of Bert4Rec and TASTE across four evaluation metrics (Recall@10/20 and NDCG@10/20).
+The results reveal a clear and consistent pattern for both backbone models.
+When $\tau$ is set too low (0.2 to 0.4), the semantic shortcut extractor becomes overly permissive, allowing many irrelevant or noisy items into the shortcut, which results in degraded performance.
+As $\tau$ increases to a moderate range (0.5 to 0.6), performance improves markedly, indicating that the shortcut extractor is now able to focus on truly relevant interactions and effectively filter out noise.
+For example, Bert4Rec achieves its best R@10 around $\tau=0.6$, and similar patterns are observed for all metrics and for the TASTE backbone.
+Importantly, our LISRec framework consistently outperforms baseline models without semantic filtering across the entire range of $\tau$ values.
+This demonstrates the robustness and generalizability of our approach–even when the threshold is not optimally tuned, shortcut extraction leads to reliable improvements in recommendation quality.
+(a) Raw User-Item Interaction Trained $M_{\text{Rec}}$.
+(b) $M_{\text{Filter}}$-Retained Items Trained $M_{\text{Rec}}$.
+Figure 4. Performance of Recommendation Models Using Different Semantic Shortcut Extraction Methods.
+5.4. Effectiveness of Different Item Representations for Shortcut Extraction
+To evaluate the effectiveness of the semantic shortcut extractor in LISRec, we perform a comparative analysis involving various item representation strategies.
+Specifically, we compare the full LISRec model with its variant without the semantic shortcut extractor (LISRec w/o Filter), as well as versions of LISRec equipped with different item representation methods within the semantic shortcut extractor.
+We implement the semantic shortcut extractor $M_{\text{Filter}}$ using three approaches: an ID-based recommendation model, Bert4Rec (Sun2019Bert4rec), and a dense retriever, BGE-Base (bge_embedding), which serve as baselines for evaluating representation effectiveness.
 (a) Average Similarity Scores among $M_{\text{Filter}}$-Retained Items.
 (b) Similarity Scores between $M_{\text{Filter}}$-Retained Items and the Target Items.
-Figure 3: Effectiveness of ConsRec in Learning Item Representations for Denoising.
-Figure 3(a) and Figure 3(b) illustrate the item similarities among $M_{\text{Filter}}$-retained items and the target items.
-(a) Raw User-Item Interaction Trained $M_{\text{Rec}}$.
-(b) $M_{\text{Filter}}$-Retained Items Trained $M_{\text{Rec}}$.
-Figure 4: Performance of Recommendation Models Using Different Item Denoising Methods.
-5.3 Effectiveness of Different Item Representation Strategies in Denoising
-To assess the effectiveness of the denoising module in ConsRec, we conduct a comparative analysis involving various item representation strategies.
-Specifically, we compare the full ConsRec model with its variant without the denoising module (ConsRec w/o Filter), as well as versions of ConsRec equipped with different item representation methods within the denoising module.
-We implement the denoising module $M_{\text{Filter}}$ using three approaches: an ID-based recommendation model, Bert4Rec Sun et al.
-(2019), and a dense retriever, BGE-Base Xiao et al.
-(2023), which serve as baselines for evaluating representation effectiveness.
+Figure 5.
+Effectiveness of LISRec in Learning Item Representations for Semantic Shortcut Extraction.
+Figure 5(a) and Figure 5(b) illustrate the item similarities among $M_{\text{Filter}}$-retained items and the ground truth items.
 As illustrated in Figure 5, we analyze the text similarity among the items retained by $M_{\text{Filter}}$ using the BGE-Base model to compute pairwise similarity scores.
-In Figure 3(a), we report the average similarity among the retained items using different denoising methods.
-The results show that all methods improve the internal similarity of the retained items.
-Notably, when using ConsRec, the retained user-interacted items achieve the highest average similarity, indicating stronger coherence among these retained items.
+In Figure 5(a), we report the average similarity among the items selected by different semantic shortcut extraction strategies.
+The results show that all methods increase the internal similarity among selected items.
+Notably, when using LISRec, the extracted item subset achieves the highest average similarity, indicating stronger semantic alignment within the shortcut.
 Furthermore, Figure 5(b) presents the average similarity between the $M_{\text{Filter}}$-retained user-interacted items and the ground-truth target item.
-Among all methods, ConsRec yields the highest similarity, demonstrating its effectiveness in retaining items that are semantically aligned with the target item that users needed.
-As illustrated in Figure 4, we evaluate the effectiveness of various denoising methods across different recommendation systems.
-Figures 4(a) and 4(b) report the performance of the recommendation module $M_{\text{Rec}}$ when trained on either the raw user-item interaction set $\mathcal{H}$ or the denoised subset $\mathcal{H}_{\text{Filter}}$ retained by $M_{\text{Filter}}$.
-In both cases, ConsRec consistently outperforms other item representation models, with particularly strong gains at larger ranking cutoffs.
-This demonstrates its effectiveness in helping the recommendation model prioritize more relevant items, thereby better satisfying user preferences.
-The advantage of ConsRec can be attributed to its PLM-based ability to comprehend item text semantics and to capture item similarity patterns from user-item interaction signals.
-After training the recommendation module $M_{\text{Rec}}$ using denoised user-item interactions, ConsRec achieves substantial performance improvements, further highlighting its effectiveness.
-5.4 The Effectiveness of Item Denoising in Enhancing the Recommendation Module
-(a) TASTE.
-(b) ConsRec.
-Figure 5: Embedding Visualization of Items in Beauty. We randomly select a user and then use t-SNE to visualize item embeddings using TASTE and ConsRec.
-As illustrated in Figure 5, we investigate how the retained items contribute to improving the recommendation module during training.
-Specifically, we get the item representations retained and removed by $M_{\text{Filter}}$ using TASTE and ConsRec.
-Figure 5(a) shows the embedding space learned by TASTE, which is trained without item denoising.
-In this setting, the model yields a homogeneous embedding space, where retained and removed items are mixed together without clear separation.
-In contrast, Figure 5(b) demonstrates that ConsRec is able to better distinguish the removed items by mapping them to the periphery of the embedding space, effectively treating them as outliers.
-This observation highlights the effectiveness of ConsRec in leveraging denoising signals as prior knowledge to guide the recommendation module.
+Among all methods, LISRec yields the highest similarity, demonstrating its effectiveness in maintaining items that are semantically aligned with the target item that users need.
+As shown in Figure 4, we further evaluate the impact of different shortcut extraction methods on overall recommendation performance.
+Figures 4(a) and 4(b) report the performance of the recommendation module $M_{\text{Rec}}$ when trained on either the raw user-item interaction set $\mathcal{H}$ or the shortcut subset $\mathcal{H}_{\text{Shortcut}}$ extracted by $M_{\text{Filter}}$.
+In both cases, LISRec consistently outperforms other item representation models, with particularly strong gains at larger ranking cutoffs.
+This demonstrates its effectiveness in enabling the recommendation model to focus on the most relevant items, thereby better capturing stable user preferences.
+The advantage of LISRec is attributed to its PLM-based ability to capture item text semantics and to identify semantically meaningful patterns in user-item interactions.
+After training the recommendation module $M_{\text{Rec}}$ using the shortcut-extracted interaction histories, LISRec achieves substantial performance improvements, further highlighting the effectiveness of semantic shortcut extraction.
+Table 5.
+Case Studies.
+We present cases from two datasets, Amazon Beauty and Yelp.
+Different colors are highlighted to distinguish preference types: Blue represents majority preference, and Orange represents discrete interaction.
+Case #1 in Amazon Beauty
+Node_ID: $n_{1}$      Name: Bare Escentuals bareMinerals Purely Nourishing Moisturizer: Combination Skin
+Description:Hydrate skin with this lightweight, remarkably effective moisturizer for improved firmness, elasticity …
+Node_ID: $n_{2}$      Name: Ole Henriksen Truth Creme Advanced Hydration, 1.7 Fluid Ounce
+Description: … An effective and groundbreaking anti-aging moisturizer contains advanced hydration complexity that deliver …
+Node_ID: $n_{3}$      Name: VALENTINO VALENTINA by Valentino for WOMEN: EAU DE PARFUM SPRAY 1.7 OZ
+Description: … Valentina By Valentino 1.7 Oz Eau De Parfum Spray
+Node_ID: $n_{4}$      Name: Boscia Oil-Free Nightly Hydration, 1.4-Fluid Ounce
+Description: … An oil-free nightly moisturizer that provides weightless hydration, combats excess oil and refines the look of …
+Similarity: $S_{1,2}=0.4441$, $S_{1,3}=0.2527$, $S_{1,4}=0.4608$, $S_{2,3}=0.2125$, $S_{2,4}=0.4477$, $S_{3,4}=0.2049$
+Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
+The maximum connected subgraph $C_{\text{max}}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Shortcut}}$ is: $\left\{1,2,4\right\}$
+Case #2 in Yelp
+Node_ID: $n_{1}$      Address: Las Vegas, NV      Name: Maggiano’s Little Italy
+Description:Nightlife, Restaurants, Italian, Bars …
+Node_ID: $n_{2}$      Address: Las Vegas, NV      Name: Lazy Dog Restaurant & Bar
+Description:Nightlife, Food, Beer, Wine & Spirits, Pizza, Burgers …
+Node_ID: $n_{3}$      Address: Las Vegas, NV      Name: Spay & Neuter Center of Southern Nevada
+Description:Veterinarians, Pets, Pet Services
+Node_ID: $n_{4}$      Address: Las Vegas, NV      Name: Weera Thai Restaurant - Sahara
+Description:Restaurants, Beer, Wine & Spirits, Food, Vegetarian, …, Seafood, Bars, Nightlife
+Similarity: $S_{1,2}=0.5926$, $S_{1,3}=0.4699$, $S_{1,4}=0.5121$, $S_{2,3}=0.4865$, $S_{2,4}=0.5150$, $S_{3,4}=0.3932$
+Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
+The maximum connected subgraph $C_{\text{max}}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Shortcut}}$ is: $\left\{1,2,4\right\}$
+(a) Similarity Scores between User Representation and Target Items.
+(b) Embedding Visualization of Items using LISRec.
+Figure 6.
+Effectiveness of Semantic Shortcut Extraction in Enhancing the Recommendation Module.
+Figure 6(a) and Figure 6(b) show the improvement of user and item representations after extracting shortcuts.
+5.5. Effectiveness of Shortcut Extraction in Enhancing the Recommendation Module
+As illustrated in Figure 6, we investigate how the shortcut-retained items contribute to improving the recommendation module during training.
+For each user, we first compute the embedding similarity between the user representation and the corresponding ground-truth item, under both the original and shortcut-applied training settings.
+Figure 6(a) shows that applying shortcut extraction during training consistently increases this similarity.
+This finding confirms that shortcut extraction strengthens the semantic alignment between user and target item representations, leading to more faithful modeling of user intent.
+Furthermore, we randomly select a user and then use t-SNE to visualize item embeddings.
+As shown in Figure 6(b), LISRec is able to better distinguish the removed items by mapping them to the periphery of the embedding space, effectively treating them as outliers.
+This observation highlights the effectiveness of LISRec in leveraging denoising signals as prior knowledge to guide the recommendation module.
 By doing so, it facilitates a more discriminative representation space, where retained and removed items are clearly separated through contrastive training.
-6 Conclusion
-This paper introduces the User-Consistent Preference based Sequential Recommendation System (ConsRec), a novel framework designed to model user-consistent preferences for denoising user-interacted items.
-By explicitly filtering out noisy items, ConsRec enhances the interpretability of denoising and demonstrates its generalization ability, extending its advantages across various recommendation models.
-Limitation
-While ConsRec demonstrates effective denoising capabilities for sequential recommendations, the approach has several limitations that must be acknowledged.
-First, the performance of ConsRec is contingent on the quality of item descriptions; incomplete or low-quality descriptions may hinder the model’s ability to effectively filter out noisy user-item interactions.
-Secondly, for cold-start users with limited interaction history, ConsRec needs to build new user-item graphs from scratch, potentially introducing latency in generating personalized recommendations.
-This limitation could be mitigated by employing incremental graph updates, which dynamically add new edges as interactions occur, thereby reducing the need for full graph recomputation.
+5.6. Case Studies
+In Table 5, we present two examples from the Amazon Beauty and Yelp datasets to evaluate the effectiveness of LISRec.
+For each dataset, we randomly select a user and showcase how semantic shortcut extraction operates on their interaction history.
+In the case studies, we detail the procedure of constructing the personalized item interaction graph and extracting the maximal preference-consistent subgraph, which serves to filter out noisy interactions.
+As shown in the first case from the Amazon Beauty dataset, the user primarily interacts with face cream products focused on moisturizing, such as “Bare Escentuals bareMinerals Purely Nourishing Moisturizer” and “Ole Henriksen Truth Creme Advanced Hydration”.
+However, the user accidentally clicks on a product–“VALENTINO VALENTINA by Valentino for WOMEN: EAU DE PARFUM SPRAY 1.7 OZ”–due to their instant interests, which are significantly different in terms of both product types and attributes.
+In detail, while other items in the sequence emphasize “hydrating” and “moisturizing”, the “Parfum Spray” is unrelated to these preferences.
+When LISRec constructs the user-interacted item graph, the “Parfum Spray” product fails to form edges with other items, indicating that this interaction is likely an outlier and should be excluded from user-interacted items.
+Consequently, the denoised item sequence $\mathcal{H}_{\text{Shortcut}}$ includes only products aligned with the consistent preferences of a specific user, such as moisturizing creams, enabling the recommendation module to better model user behavior from the user-item interactions.
+In the second case, we present a specific user from the Yelp dataset who visits various locations.
+This user primarily interacts with places characterized by attributes such as “Nightlife”, “Restaurants”, and “Bars”, demonstrating a consistent pattern of preferences.
+However, one of the interacted locations–the “Spay & Neuter Center of Southern Nevada”–is a pet hospital, which is distinct from the user-consistent preference for nightlife and dining.
+Although this pet hospital is situated on the same street as other frequently visited locations, it likely represents an occasional visit rather than a reflection of the general behavior of the user.
+LISRec assigns a lower similarity score between this noisy item and other user-interacted items, effectively excluding the noisy item from the sequence of user-interacted history.
+This enables the recommendation model to focus on more representative items, such as “Maggiano’s Little Italy” and “Weera Thai Restaurant”.
+6. Conclusion
+This paper introduces the Learned Item Shortcuts for Recommendation System (LISRec), a new framework that models stable user preferences by identifying dominant item subsets within a latent semantic space.
+LISRec leverages universal item representations to construct semantically meaningful shortcuts in user interaction histories.
+Our experiments on Yelp and Amazon Product datasets show that LISRec consistently outperforms baseline and previous text-based models.
+The results demonstrate that instruction-tuned universal item representations enable more robust shortcut selection and lead to more accurate, preference-consistent recommendations.
+Acknowledgements.
+This work is partly supported by the National Natural Science Foundation of China (No. 62206042 and No. 62461146205). This work is also supported by the AI9Stars community.
 Appendix A Appendix
-A.1 License
+A.1. License
 We show the licenses of the datasets that we use.
-Yelp [^2] https://www.yelp.com/dataset uses Apache License 2.0, while Amazon Product shows its terms of use at website [^3] http://jmcauley.ucsd.edu/data/amazon/.
+Yelp [^5] https://www.yelp.com/dataset uses Apache License 2.0, while Amazon Product shows its terms of use on website [^6] http://jmcauley.ucsd.edu/data/amazon/.
 All of these licenses and agreements allow their data for academic use.
-A.2 Additional Experimental Details of Data Preprocessing
+A.2. Additional Experimental Details of Data Preprocessing
+In our experiments, we employ Recbole (zhao2021recbole) for preprocessing the data for sequential recommendation modeling, maintaining the same experimental settings as prior works (Zhou2020s3; Xie2022DIF-SR; liu2023text; chen2024hllm).
+Specifically, each user-item interaction sequence $\{v_{1},\dots,v_{t}\}$ is processed using the leave-one-out strategy (Xie2022DIF-SR; chen2018sequential; Sun2019Bert4rec; Zhou2020s3) to construct the training, development, and testing sets.
+In detail, the testing set predicts the $t$-th item $v_{t}$ based on the previously interacted items $v_{1:t-1}$ of the user, while the development set predicts the t1t-1-th item $v_{t-1}$ given the interactions $v_{1:t-2}$.
+The training set predicts $v_{i}$ using $v_{1:i-1}$, where 1<i<t11<i<t-1.
+Items and users with fewer than five interactions are filtered out, and all user-item interactions are treated as implicit feedback (chen2018sequential; Sun2019Bert4rec; Zhou2020s3; Xie2022DIF-SR; liu2023text; chen2024hllm).
+The data statistics are shown in Table 6.
+Table 6. Statistics of the Datasets Used for Modeling the Sequential Recommendation Module ($M_{\text{Rec}}$).
 Dataset Data Information Split
@@
 Yelp 30,499 20,068 317,182 225,685 30,499 30,499
-Table 4: Statistics of the Datasets Used for Modeling the Sequential Recommendation Module ($M_{\text{Rec}}$).
+Table 7. Statistics of Training Data Used in the Filter Module ($M_{\text{Filter}}$). The data is used for optimizing item representations to estimate their similarities.
 Dataset Data Information Split
@@
 Total 107,602 62,450 979,618 902,740 77,648
-Table 5: Statistics of Training Data Used in the Filter Module ($M_{\text{Filter}}$). The data is used for optimizing item representations to estimate their similarities.
-In our experiments, we employ Recbole Zhao et al.
-(2021) for preprocessing the data for sequential recommendation modeling, maintaining the same experimental settings as prior works Zhou et al.
-(2020); Xie et al.
-(2022); Liu et al.
-(2023); Chen et al.
-(2024).
-Specifically, each user-item interaction sequence $\{v_{1},\dots,v_{t}\}$ is processed using the leave-one-out strategy Xie et al.
-(2022); Chen et al.
-(2018); Sun et al.
-(2019); Zhou et al.
-(2020) to construct the training, development, and testing sets.
-In detail, the testing set predicts the $t$-th item $v_{t}$ based on the previously interacted items $v_{1:t-1}$ of the user, while the development set predicts the $t-1$-th item $v_{t-1}$ given the interactions $v_{1:t-2}$.
-The training set predicts $v_{i}$ using $v_{1:i-1}$, where $11<i<t-$ 1.
-Items and users with fewer than five interactions are filtered out, and all user-item interactions are treated as implicit feedback Chen et al.
-(2018); Sun et al.
-(2019); Zhou et al.
-(2020); Xie et al.
-(2022); Liu et al.
-(2023); Chen et al.
-(2024).
-The data statistics are shown in Table 4.
 Then, we construct the dataset for training the language model to represent items for estimating their similarities (Eq.
-11).
+10).
 We start with the training and development sets used in sequential recommendation modeling and further process them for the Next Item Prediction (NIP) task.
@@
 Then, we balance the number of data points to construct the training and development sets.
-This approach mitigates the risk of overfitting caused by data imbalance of different recommendation tasks.
+This approach mitigates the risk of overfitting caused by data imbalance in different recommendation tasks.
 As shown in Table 7, the data statistics for both the training and development sets used in item representation learning are presented.
-Following previous studies Xie et al.
-(2022); Liu et al.
-(2023), we conduct evaluations in the full ranking testing scenario Dallmann et al.
-(2021); Krichene and Rendle (2020), where items are ranked across the entire item set rather than sampling subsets.
+Following previous studies (Xie2022DIF-SR; liu2023text), we conduct evaluations in the full ranking testing scenario (dallmann2021case; krichene2022sampled), where items are ranked across the entire item set rather than sampling subsets.
 This evaluation set mitigates the inconsistencies often observed in sampled evaluations, offering a more realistic and comprehensive assessment of recommendation performance.
-A.3 Additional Details of Baseline Models
-We reproduce all item-ID based models and DWSRec Zhang et al.
-(2024b) using implementations provided by RecBole Zhao et al.
-(2021), maintaining consistent settings across all models for fair comparison.
-For CoWPiRec and TASTE, we adhere to the default experimental settings as specified in their original papers.
-For HLLM, due to computational constraints, we ultilize TinyLlama-1.1B Zhang et al.
-(2024c) as the item-LLM component, and the user-LLM is implemented as a custom Llama model Touvron et al.
-(2023) with two hidden layers, trained from scratch using the following settings: learning rate of 1e-4, batch size of 2, training for 5 epochs, and a negative sampling ratio of 1024 negative samples per positive sample.
-(a) SASRec
-(b) BERT4Rec
-(c) TASTE
-Figure 6: Hyperparameter Sensitivity Analysis.
-We conduct experiments on the threshold hyperparameter, $\tau$, with results derived from the Amazon Beauty dataset using the SASRec 6(a), BERT4Rec 6(b), and TASTE 6(c) backbone models.
-A.4 Hyperparameter Sensitivity Analysis
-We investigate the sensitivity of the key hyperparameter $\tau$, which determines the similarity threshold during user-item interaction graph construction, as described in Sec.
-3.2.
-We systematically vary $\tau$ from 0.2 to 0.8 while maintaining all other experimental settings.
-Figure 6 illustrates the performance across four evaluation metrics (Recall@10/20 and NDCG@10/20) for three backbone models: SASRec, Bert4Rec, and TASTE.
-The results reveal a consistent pattern across all models and metrics.
-At lower $\tau$ values (0.2-0.4), the filtering mechanism is overly permissive, allowing to much noise in the interaction data and resulting in suboptimal performance.
-As $\tau$ increases to moderate values (0.5-0.6), we observe substantial performance improvements as irrelevant interactions are effectively filtered out.
-For instance, SASRec achieves peak R@10 performance at approximately $\tau=0.6$, with similar bell-shaped trends evident across other models and metrics.
-When $\tau$ exceeds 0.7, the filtering becomes excessively strict, potentially discarding valuable interaction signals and causing performance deterioration.
-A.5 Ablation Study on Different Filter Models
-To evaluate the effectiveness of the $M_{\text{Filter}}$ component in ConsRec, we conduct ablation studies comparing our approach with two alternative filter models–a vanilla T5 $M_{\text{Filter}}$ initialized from the T5-base checkpoint without additional training, and a TASTE $M_{\text{Filter}}$ trained specifically on recommendation data.
-Dataset Method R@20 N@20
-Beauty ConsRec 0.1774 0.0772
-w/ T5 $M_{\text{Filter}}$ 0.1761 0.0762
-w/ TASTE $M_{\text{Filter}}$ 0.1767 0.0767
-Sports ConsRec 0.1082 0.0496
-w/ T5 $M_{\text{Filter}}$ 0.1073 0.0457
-w/ TASTE $M_{\text{Filter}}$ 0.1077 0.0471
-Toys ConsRec 0.1896 0.0826
-w/ T5 $M_{\text{Filter}}$ 0.1879 0.0802
-w/ TASTE $M_{\text{Filter}}$ 0.1883 0.0811
-Yelp ConsRec 0.1173 0.0509
-w/ T5 $M_{\text{Filter}}$ 0.1169 0.0507
-w/ TASTE $M_{\text{Filter}}$ 0.1170 0.0508
-Table 6: Ablation Study on Different Filter Models in ConsRec.
-As shown in Table 6, ConsRec consistently outperforms both alternative implementations across all datasets, demonstrating the effectiveness of our proposed training strategy for the filter module.
-The vanilla T5 $M_{\text{Filter}}$ exhibits the weakest performance, as it lacks domain-specific knowledge for recommendation tasks.
-The TASTE $M_{\text{Filter}}$ shows improved performance due to its training on recommendation tasks, but its effectiveness is limited by exposure to only a single dataset during training.
-In contrast, our ConsRec $M_{\text{Filter}}$ benefits from data-balanced pre-training across multiple datasets, resulting in more robust and generalizable item representations.
-This confirms that our training approach effectively enhances the model’s ability to capture relevant item characteristics while maintaining strong generalization capabilities across different recommendation scenarios.
-Case #1 in Amazon Beauty
-Node_ID: $n_{1}$    Name: Bare Escentuals bareMinerals Purely Nourishing Moisturizer: Combination Skin
-Description:Hydrate skin with this lightweight, remarkably effective moisturizer for improved firmness, elasticity …
-[…]
-Description: … An effective and groundbreaking anti-aging moisturizer contains advanced hydration complexity that deliver …
-[…]
-Description: … Valentina By Valentino 1.7 Oz Eau De Parfum Spray
-[…]
-Description: … An oil-free nightly moisturizer that provides weightless hydration, combats excess oil and refines the look of …
-Similarity: $S_{1,2}=0.4441$, $S_{1,3}=0.2527$, $S_{1,4}=0.4608$, $S_{2,3}=0.2125$, $S_{2,4}=0.4477$, $S_{3,4}=0.2049$
-Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
-The maximum connected subgraph $C_{max}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Filter}}$ is: $\left\{1,2,4\right\}$
-Case #2 in Yelp
-Node_ID: $n_{1}$    Address: Las Vegas, NV    Name: Maggiano’s Little Italy
-Description:Nightlife, Restaurants, Italian, Bars …
-[…]
-Description:Nightlife, Food, Beer, Wine & Spirits, Pizza, Burgers …
-[…]
-Description:Veterinarians, Pets, Pet Services
-[…]
-Description:Restaurants, Beer, Wine & Spirits, Food, Vegetarian, …, Seafood, Bars, Nightlife
-Similarity: $S_{1,2}=0.5926$, $S_{1,3}=0.4699$, $S_{1,4}=0.5121$, $S_{2,3}=0.4865$, $S_{2,4}=0.5150$, $S_{3,4}=0.3932$
-Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
-The maximum connected subgraph $C_{max}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Filter}}$ is: $\left\{1,2,4\right\}$
-Table 7: Case Studies.
-We present cases from two datasets, Amazon Beauty and Yelp.
-Different colors are highlighted to distinguish preference types: Blue represents majority preference, and Orange represents discrete interaction.
-A.6 Case Studies
-In Table 7, we present two examples from the Amazon Beauty and Yelp datasets to evaluate the effectiveness of ConsRec.
-For each dataset, a user is randomly selected, and we illustrate the process of denoising user-interacted items by modeling user-consistent preferences.
-In the case studies, we detail the procedure of identifying the maximum connected subgraph from the user-interacted item graph for item denoising.
-As shown in the first case from the Amazon Beauty dataset, the user primarily interacts with face cream products focused on moisturizing, such as “Bare Escentuals bareMinerals Purely Nourishing Moisturizer” and “Ole Henriksen Truth Creme Advanced Hydration”.
-However, the user accidentally clicks on a product–“VALENTINO VALENTINA by Valentino for WOMEN: EAU DE PARFUM SPRAY 1.7 OZ”–due to their instant interests, which are significantly different in terms of both product types and attributes.
-In detail, while other items in the sequence emphasize “hydrating” and “moisturizing”, the “Parfum Spray” is unrelated to these preferences.
-When ConsRec constructs the user-interacted item graph, the “Parfum Spray” product fails to form edges with other items, indicating that this interaction is likely an outlier and should be excluded from user-interacted items.
-Consequently, the denoised item sequence $\mathcal{H}_{\text{Filter}}$ includes only products aligned with the consistent preferences of a specific user, such as moisturizing creams, enabling the recommendation module to better model user behavior from the user-item interactions.
-In the second case, we present a specific user from the Yelp dataset who visits various locations.
-This user primarily interacts with places characterized by attributes such as “Nightlife”, “Restaurants”, and “Bars”, demonstrating a consistent pattern of preferences.
-However, one of the interacted locations– the “Spay & Neuter Center of Southern Nevada”–is a pet hospital, which is distinct from the user-consistent preference for nightlife and dining.
-Although this pet hospital is situated on the same street as other frequently visited locations, it likely represents an occasional visit rather than a reflection of the general behavior of the user.
-ConsRec assigns a lower similarity score between this noisy item and other user-interacted items, effectively excluding the noisy item from the sequence of user-interacted history.
-This enables the recommendation model to focus on more representative items, such as “Maggiano’s Little Italy” and “Weera Thai Restaurant”.
-A.7 Theoretical Analysis of Consistent User Preference Modeling via ConsRec
-Preliminary. Let =$\mathcal{H}=\{v_{1},\dots,v_{n}\}$ be the full interaction history of a user. The consistent user preference set $P$ can be represented:
-$P\subseteq\mathcal{H},\lvert P\rvert=m,$ (17)
+A.3. Theoretical Analysis of Stable User Preference Modeling via LISRec
+Preliminary. Let $\mathcal{H}=\{v_{1},\dots,v_{n}\}$ be the full interaction history of a user. The consistent user preference set $P$ can be represented:
+(15) $\small P\subseteq\mathcal{H},\lvert P\rvert=m,$
 and the remaining items set can be denoted:
-$N=\mathcal{H}\setminus P,\lvert N\rvert=n-m.$ (18)
+(16) $\small N=\mathcal{H}\setminus P,\lvert N\rvert=n-m.$
 Then, we denote the user interaction graph as an undirected graph:
-$G_{\mathcal{U}}=(\mathcal{H},\mathcal{E}).$ (19)
+(17) $\small G_{\mathcal{U}}=(\mathcal{H},\mathcal{E}).$
 Here $(v_{i},v_{j})\in\mathcal{E}\Longleftrightarrow\cos(\vec{v}_{i},\vec{v}_{j}\bigr)=S_{ij}\geq\tau$, where the fixed threshold $\tau\in(0,1)$.
@@
 We denote the collection of all connected components as:
-$\mathcal{C}=\{C_{1},C_{2},\dots,C_{\ell}\},$ (20)
+(18) $\small\mathcal{C}=\{C_{1},C_{2},\dots,C_{\ell}\},$
 where each connected component $C_{k}\subseteq\mathcal{H}$ and $C_{i}\cap C_{j}=\emptyset$ for all $i\neq j$. Then, we denote
-$C_{\text{max}}=\mathrm{arg}\max_{1\leq k\leq\ell}\lvert C_{k}\rvert$ (21)
+(19) $\small C_{\text{max}}=\mathrm{arg}\max_{1\leq k\leq\ell}\lvert C_{k}\rvert$
 as the connected component with the maximum number of nodes. Above all, our main claim is that the maximum connected component $C_{\text{max}}$ is exactly the user preference set $P$.
 Assumptions. We need the following three weakest conditions to support the subsequent derivation.
-(i) Every pair of nodes belongs to $P$ is similar:
-$\forall v_{i},v_{j}\in P:S_{ij}\geq\tau.$ (22)
+(i) The preference nodes are mutually reachable in $G_{\mathcal{U}}$:
+(20) $\small\forall v_{i},v_{j}\in P:v_{i}\sim_{\tau}v_{j}$
 (ii) There is no edge between any noise node in $N$ and any preferred node in $P$:
-$\forall v_{i}\in P,v_{j}\in N:S_{ij}<\tau;$ (23)
+(21) $\small\forall v_{i}\in P,v_{j}\in N:S_{ij}<\tau;$
 (iii) Even if noise nodes have high similarity and form small clusters, their scale will not exceed that of the preference set $P$:
-$\forall\{C_{1}^{N},\ldots,C_{q}^{N}\}\subseteq\mathcal{C}:\max_{1\leq r\leq q}\lvert C_{r}^{N}\rvert\mathrm{<}m.$ (24)
-Lemmas. The following three lemmas are correspond to assumptions (i)-(iii), respectively.
-(1) The preference set is the connected component.
-By assumption (i), $\forall v_{i},v_{j}\in P$, then $S_{ij}\geq\tau$, so that $(v_{i},v_{j})\in\mathcal{E}$.
-If any two nodes in the set $P$ are directly connected to each other, then $P$ forms a complete graph, which necessarily constitutes a connected component.
+(22) $\small\forall\{C_{1}^{N},\ldots,C_{q}^{N}\}\subseteq\mathcal{C}:\max_{1\leq r\leq q}\lvert C_{r}^{N}\rvert\mathrm{<}m.$
+Lemmas. The following three lemmas correspond to assumptions (i)-(iii), respectively.
+(1) The preference set induces a connected component.
+By assumption (i), every pair of nodes in $P$ can reach each other through edges whose similarities exceed $\tau$, hence all nodes of $P$ lie in the same connected component of $G_{\mathcal{U}}$.
+Moreover, by assumption (ii), no node in $N$ is adjacent to any node in $P$, so that component contains only the nodes of $P$.
 (2) Noise nodes and preference nodes are not in the same component.